### v1

# Homogeneous $\mathrm{G}_{2}$ and Sasakian instantons on the Stiefel $7$-manifold

Abstract
We study homogeneous instantons on the seven dimensional Stiefel manifold $V^{5,2}$ in the context of $\mathrm{G}_{2}$ and Sasakian geometry. According to the reductive decomposition of $V^{5,2}$, we provide an explicit description of all invariant $\mathrm{G}_{2}$ and Sasakian structures. In particular, we characterise the invariant $\mathrm{G}_{2}$-structures inducing a Sasakian metric, among which the well known nearly parallel $\mathrm{G}_{2}$-structure (Sasaki-Einstein) is included. As a consequence, we classify the invariant connections on homogeneous principal bundles over $V^{5,2}$ with gauge group ${\rm U}(1)$ and ${\rm SO}(3)$, satisfying either the $\mathrm{G}_{2}$ or the Sasakian instanton condition. Finally, we analyse the Yang Mills condition for those invariant connections.
1 Introduction
The Stiefel manifold $V^{n,k}$ of orthonormal $k$-frames in the $n$-dimensional Euclidean space (i.e. $1\leq k\leq n$) admits a transitive ${\rm SO}(n)$-action with isotropy subgroup isomorphic to ${\rm SO}(n-k)$, thus, it can be written homogeneously as $V^{n,k}={\rm SO}(n)/{\rm SO}(n-k)$ and consequently, its family of ${\rm SO}(n)$-invariant Riemannian metrics are described in terms of the isotropy representation of $V^{n,k}$. For instance, for $k=1$ or $k=n$, the isotropy representation is irreducible (indeed $V^{n,1}=S^{n}$ or $V^{n,n}={\rm SO}(n)$), in those cases, there is a unique ${\rm SO}(n)$-invariant metric, which is Einstein [wolf1968]. For $k=2$, the isotropy representation is reducible and the space of ${\rm SO}(n)$-invariant Riemannian metrics is described by a $4$-dimensional family, within there is a unique Einstein metric, which is inherited from the Grasmmanian $G_{2}(\mathbb{R}^{n})$ (e.g. [Arvanitoyeorgos1996, kerr1998new]). In addition, $V^{n,2}$ fits as an instance of the Boothby-Wang construction, as a total space of a $S^{1}$-principal bundle [boothby1958]*Theorem 3, moreover, it has a regular contact structure whose associated Reeb vector field generates a $1$-parameter subgroup of right translations of $V^{n,2}$.
A contact structure [^1] The reader unfamiliar with Sasakian geometry can find a complete study on this subject in [Boyer2008] on a $n$-dimensional manifold $M$ is a 1-form satisfying $\eta\wedge(d\eta)^{n}\neq 0$, in this case there exist a unique vector $\xi\in\mathscr{X}(M)$, called Reeb vector field, such that $\eta(\xi)=1$. The contact structure is called an almost contact structure if there exist an endomorphism $\Phi\in{\mathrm{End}}(TM)$ which satisfy $\Phi^{2}=-{\rm I}_{TM}+\eta\otimes\xi.$ A Riemannian metric $g$ is called compatible with the contact structure, if $g$ satisfies $g(\Phi X,\Phi Y)=g(X,Y)-\eta(X)\eta(Y),$ in this case, the contact structure is called an almost contact metric structure. In addition, if $\mathcal{S}=(\eta,\xi,g,\Phi)$ satisfies $2g(X,\Phi(Y))=d\eta(X,Y)$, the quadruple $\mathcal{S}$ is called a contact metric structure. Furthermore, if the Nijenhuis tensor $N_{\Phi}$ induced by $\Phi$ satisfies $N_{\Phi}=-d\eta\otimes\xi$, then $\mathcal{S}$ is called a Sasakian structure. In this paper, we consider the $7$–dimensional Stiefel manifold $V^{5,2}$ endowed with a one-parameter family of ${\rm SO}(5)$-invariant Sasakian structures $\mathcal{S}=(\eta,\xi,g,\Phi)$. For instance, homogeneous contact/Sasakian structures have been study in [andrada2009, correa2019] and [perrone1998homogeneous].
On the other hand, a differential $3$-form $\varphi$ on an oriented $7$–dimensional manifold $M$ is called a $\mathrm{G}_{2}$–structure if at each $p\in M$, there is a basis $e_{1},\dots,e_{7}$ of $T_{p}M$ such that
$\varphi_{p}=e^{125}+e^{136}+e^{147}+e^{234}-e^{267}+e^{357}-e^{456},$ (1.1)
where $e^{1},\dots,e^{7}$ denotes the dual basis of $T_{p}^{\ast}M$ and $e^{ijk}=e^{i}\wedge e^{j}\wedge e^{k}$. In particular, $\varphi$ determines a Riemannian metric $g_{\varphi}$, we study the family of ${\rm SO}(5)$-invariant $\mathrm{G}_{2}$-structures on $V^{5,2}$, such that the induced metric $g_{\varphi}$ is Sasakian. With respect to these structures, we study the theoretical Gauge instanton equations associated with the underlying Sasakian/$\mathrm{G}_{2}$-geometry of $V^{5,2}$. Namely, on a $G$–principal bundle $P\to M^{7}$ with $\sigma\in\Omega^{3}(M)$, the $1$-form connection $A\in\mathcal{A}(P)$ is called an $\lambda$-instanton, if it satisfies (e.g. [donaldson1998gauge, Tian2000])
$\ast(\sigma\wedge F_{A})=\lambda F_{A},\quad\text{where}\quad\lambda\in\mathbb{R}$ (1.2)
where $F_{A}=dA+A\wedge A$ denotes the corresponding curvature $2$-form. According to the specific pair $(\sigma,\lambda)$ in (1.2), we define:
  1. 1.
    A Self-dual contact Instantons (SDCI) [portilla2023] is a solutions of (1.2) with $\lambda=1$ and $\sigma:=\eta\wedge\omega$ naturally defined by a contact structure of $(M,\eta)$, where $\omega:=\frac{1}{2}d\eta$ is called the fundamental $2$-form..
  2. 2.
    A $\mathrm{G}_{2}$–instantons [corrigan1983first, donaldson1998gauge] is a solutions of (1.2) with $\lambda=-1$ and $\sigma=\varphi$ a coclosed (i.e. $d\ast\varphi=0$) $\mathrm{G}_{2}$–structure on $M$.
Furthermore, for $E\to(M^{7},\mathcal{S})$ a Sasakian holomorphic bundle [biswas2010vector], the usual notion of Hermitian Yang-Mills (HYM) connection on a Kähler manifolds is extended to odd dimension in the natural fashion to the so-called transverse Hermitian Yang-Mills connections (tHYM) (see (3.1)). The concept of Chern connections also extend naturally to a Sasakian version, as a connection mutually compatible with the transverse holomorphic structure and the Hermitian metric. A noteworthy result is that if $E\to M$ is a holomorphic Sasakian bundle on a contact Calabi-Yau manifold $M$, along compatible connections the notion of SDCI, $\mathrm{G}_{2}$-instantons and tHYM coincide. cf. [portilla2023, Teorem 1.1] and [calvo2020gauge, Lemma  21].
In [donaldson1998gauge, Donaldson2011], Donaldson, Segal and Thomas suggested that it may be possible to use $\mathrm{G}_{2}$-instantons to construct an enumerative invariant of $\mathrm{G}_{2}$-manifolds. For instance, when $(M^{7},\varphi)$ has a torsion-free $\mathrm{G}_{2}$-structure $\varphi$ (i.e. $d\varphi=0$ and $d\ast\varphi=0$ [fernandez1982riemannian]), solutions to the $\mathrm{G}_{2}$-instanton equation have been studied in [sa2015g2, walpuski2015g_2, clarke2014instantons, lotay20182]. However, the construction of metrics with holonomy $\mathrm{G}_{2}$ has proven to be a challenging task, thus, the torsion-free condition has been relaxed to the class of coclosed $\mathrm{G}_{2}$-structures [ball2019gauge, clarke2023], which has a notably application in theoretical physics to the Killing spinor equations in supergravity (cf. [Ivanov2005, Ivanov10, delaOssaG21, Lotay2022, Fernandez2011]). Analogously, Wang has proposed studying the compactification of the moduli space of contact instantons as a tool to construct a Donaldson-type invariant for contact $7$-manifolds [wang2012higher].
Description of the main results: Let $\mathfrak{so}(5)=\mathfrak{so}(3)\oplus{\mathfrak{m}}$ be a reductive decomposition of $V^{5,2}$, in [reidegeld2010spaces] Reidegeld proved that the space of ${\rm SO}(5)$-invariant $\mathrm{G}_{2}$-structures has dimension $5$ and the space of induced ${\rm SO}(5)$-invariant Riemannian metrics has dimension $4$. Thus, we prove in Lemma 2.3 that the $\mathrm{Ad}({\rm SO}(3))$-invariant $3$-form (2.11) determined by $(a,b,x,y,z)\in\mathbb{R}^{5}$ defines a $\mathrm{G}_{2}$-structure (up to $\mathrm{G}_{2}$-transformation), if
$a^{2}+by>0,\quad b^{2}+ax>0\quad z\neq 0\quad\text{and}\quad xy=ab.$
In particular, it defines a family of (non-purely) coclosed $\mathrm{G}_{2}$-structures, within the nearly parallel $\mathrm{G}_{2}$-structures arises as an $S^{1}$-family for $x=a$, $y=b$ and $a^{2}+b^{2}=z^{2}=\frac{27}{512}$ (cf. Theorem 2.1).
In addition, for any $y_{1},y_{2},y_{3}\in\mathbb{R}^{+}$, Theorem 2.2 characterises the ${\rm SO}(5)$-invariant Sasakian metrics as
$y_{1}=4y_{2}^{2}\quad\text{and}\quad y_{2}=y_{3},$
as a consequence, for a given $(z,a,b,x,y)\in\mathbb{R}^{5}$, we describe the family of ${\rm SO}(5)$-invariant $\mathrm{G}_{2}$-structures inducing a Sasakian metric (cf. Corollary 2.8).
In Section 3, we revisit the interaction between SDCI, $\mathrm{G}_{2}$-instantons and tHYM connections addressed in [portilla2023, Proposition 3.11] and [calvo2020gauge, Lemma 21] against the backdrop of contact Calabi-Yau manifolds. For a bundle $E\to V^{5,2}$, in Lemma 3.1 we prove that if the connection $A$ is SDCI then it is a $\mathrm{G}_{2}$-instanton. Conversely, for a holomorphic Sasakian bundle $E\to V^{5,2}$, if the Chern connection is a $\mathrm{G}_{2}$-instanton, then it is a SCDI. Finally, we prove that HYM connections on a holomorphic bundle $E\to G_{2}(\mathbb{R}^{5})$ on the Grassmanian, pulling back to $\mathrm{G}_{2}$-instantons on $\pi^{\ast}E\to V^{5,2}$ (cf. Corollary 3.4). We recall here that the key argument follow just in presence of a SU$(3)$-structure.
Finally, in Section 4 we provide examples of homogeneous SCDI and $\mathrm{G}_{2}$-instantons over $V^{5,2}$, with gauge group ${\rm SO}(3)$ and ${\rm U}(1)$. In particular, we prove that the ${\rm SO}(3)$-invariant $\mathrm{G}_{2}$-instanton (4.8) is a Yang-Mills connection on $(V^{5,2},\varphi)$, for any ${\rm SO}(5)$-invariant coclosed $\mathrm{G}_{2}$-structure (cf. Theorem 4.1), moreover, it describes a local minimizer of the Yang-Mills functional (cf. Corollary 4.11).
Outline: In Section 2, we provide some preliminaries about $V^{5,2}$ as a homogeneous space, with the aim of characterising all the ${\rm SO}(5)$-invariant coclosed $\mathrm{G}_{2}$-structures and the associated Sasakian metrics. In Section 3, we explore the interaction between SDCI and $\mathrm{G}_{2}$–instantons on $V^{5,2}$, in particular, Lemmata 3.1, 3.3 and Corollary 3.4 indicate conditions such that contact instantons and $\mathrm{G}_{2}$-instantons coincides. Finally, in Section 4 we consider homogeneous principal bundles on $V^{5,2}$ and we construct examples of invariant instantons on ${\rm SO}(3)$, $U(1)$–homogeneous bundles. In particular, in Section 4.1 we characterise all invariant ${\rm SO}(3)$–connection and we show that these satisfy the $\mathrm{G}_{2}$ and SDCI condition for appropriate parameters and we analyse its stability as Yang-Mills connections.
Acknowledgments and funding: AM was funded by the São Paulo Research Foundation (Fapesp) [2021/08026-5] and [2023/13780-6]. LP was funded by Centre Henri Lebesgue (CHL) under the grant ANR-11-LABX-0020-01, he also would like to thanks to LMBA at Université de Bretagne Occidentale UBO for the hospitality.
2 Homogeneous $\mathrm{G}_{2}$ and contact metric structures on $V^{5,2}$
In this section, we provide an explicit description of a $\mathrm{Ad}({\rm SO}(3))$-invariant $\mathrm{G}_{2}$-structure and the induced inner product on $V^{5,2}$. As well as, we establish the notation that we will use throughout this work.
2.1 $\mathrm{G}_{2}$-structures on $V^{5,2}$
In order to describe the space of ${\rm SO}(5)$-invariant $\mathrm{G}_{2}$-structure on $V^{5,2}={\rm SO}(5)/{\rm SO}(3)$ [^2] This subject is brought up in [friedrich1997nearly]., we fix a reductive decomposition of $V^{5,2}$ as a homogeneous space. Consider the Lie algebra of ${\rm SO}(5)$
$\mathfrak{so}(5):=\{A\in\mathfrak{gl}(\mathbb{R}^{5})\;|\;A+A^{T}=0\},$ (2.1)
and denote by $E_{ij}$, the $5\times 5$ matrix with $1$ at the $(ij)$-entry and 00 elsewhere, and fix a basis for $\mathfrak{so}(5)$:
$\begin{split}e_{1}=&E_{12}-E_{21},\quad e_{4}=E_{15}-E_{51},\quad e_{7}=E_{25}-E_{52},\quad e_{10}=E_{54}-E_{45},\\ e_{2}=&E_{13}-E_{31},\quad e_{5}=E_{23}-E_{32},\quad e_{8}=E_{34}-E_{43},\quad\\ e_{3}=&E_{14}-E_{41},\quad e_{6}=E_{24}-E_{42},\quad e_{9}=E_{35}-E_{53}.\end{split}$ (2.2)
We consider the Lie algebra $\mathfrak{so}(3)$ embedded in $\mathfrak{so}(5)$ and it is spanned by $\{e_{8},e_{9},e_{10}\}$. The Lie bracket of $\mathfrak{so}(5)$ is given by the commutator describe in Table 1.
[,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] $e_{1}$ $e_{2}$ $e_{3}$ $e_{4}$ $e_{5}$ $e_{6}$ $e_{7}$ $e_{8}$ $e_{9}$ $e_{10}$
$e_{1}$ 00 $-e_{5}$ $-e_{6}$ $-e_{7}$ $e_{2}$ $e_{3}$ $e_{4}$ 00 00 00
$e_{2}$ $e_{5}$ 00 $-e_{8}$ $-e_{9}$ $-e_{1}$ 00 00 $e_{3}$ $e_{4}$ 00
$e_{3}$ $e_{6}$ $e_{8}$ 00 $-e_{10}$ 00 $-e_{1}$ 00 $-e_{2}$ 00 $e_{4}$
$e_{4}$ $e_{7}$ $e_{9}$ $e_{10}$ 00 00 00 $-e_{1}$ 00 $-e_{2}$ $-e_{3}$
$e_{5}$ $-e_{2}$ $e_{1}$ 00 00 00 $-e_{8}$ $-e_{9}$ $e_{6}$ $e_{7}$ 00
$e_{6}$ $-e_{3}$ 00 $e_{1}$ 00 $e_{8}$ 00 $-e_{10}$ $-e_{5}$ 00 $e_{7}$
$e_{7}$ $-e_{4}$ 00 00 $e_{1}$ $e_{9}$ $e_{10}$ 00 00 $-e_{5}$ $-e_{6}$
$e_{8}$ 00 $-e_{3}$ $e_{2}$ 00 $-e_{6}$ $e_{5}$ 00 00 $-e_{10}$ $e_{9}$
$e_{9}$ 00 $-e_{4}$ 00 $e_{2}$ $-e_{7}$ 00 $e_{5}$ $e_{10}$ 00 $-e_{8}$
$e_{10}$ 00 00 $-e_{4}$ $e_{3}$ 00 $-e_{7}$ $e_{6}$ $-e_{9}$ $e_{8}$ 00
Table 1: Bracket of $\mathfrak{so}(5)$ with respect to the basis (2.2)
On $\mathfrak{so}(5)$, we fix the inner product $\langle a,b\rangle=Q(a,b)=\frac{1}{2}\mathop{\mathrm{Tr}}\nolimits(ab^{T}),$ with respect to which, we have the orthogonal splitting
$\mathfrak{so}(5)={\mathfrak{m}}\oplus\mathfrak{so}(3)\quad\text{where}\quad{\mathfrak{m}}:=\mathfrak{so}(3)^{\perp}=\operatorname{{\rm Span}}\{e_{1},...,e_{7}\}.$ (2.3)
Note that in terms of the basis (2.2), the restriction of $\mathrm{ad}(e_{k})_{{\mathfrak{m}}}$ to ${\mathfrak{m}}$ are given by
$\mathrm{ad}(e_{8})_{\mathfrak{m}}$ $=\mathrm{diag}(0,A_{1},A_{1}),$
$\mathrm{ad}(e_{9})_{\mathfrak{m}}$ $=\mathrm{diag}(0,A_{2},A_{2}),$
$\mathrm{ad}(e_{10})_{\mathfrak{m}}$ $=\mathrm{diag}(0,A_{3}A_{3}),$
where $A_{1}=\begin{pmatrix}0&-1&0\\ 1&0&0\\ 0&0&0\end{pmatrix}$,  $A_{2}=\begin{pmatrix}0&0&-1\\ 0&0&0\\ 1&0&0\end{pmatrix}$ and $A_{3}=\begin{pmatrix}0&0&0\\ 0&0&-1\\ 0&1&0\end{pmatrix}$ are the canonical basis of $\mathfrak{so}(3)$. Denote by ${\mathfrak{m}}_{1}=\operatorname{{\rm Span}}\{e_{1}\},$ ${\mathfrak{m}}_{2}:=\operatorname{{\rm Span}}\{e_{2},e_{3},e_{4}\}$ and ${\mathfrak{m}}_{3}:=\operatorname{{\rm Span}}\{e_{5},e_{6},e_{7}\}.$ By using the Lie bracket in Table 1 is straightforward to verify that
$\mathrm{ad}(\mathfrak{so}(3)){\mathfrak{m}}_{1}=0,\quad\mathrm{ad}(\mathfrak{so}(3)){\mathfrak{m}}_{2}\subseteq{\mathfrak{m}}_{2}\quad\text{and}\quad\mathrm{ad}(\mathfrak{so}(3)){\mathfrak{m}}_{3}\subseteq{\mathfrak{m}}_{3}.$
Consequently, for the reductive decomposition (2.3) we obtain the irreducible $\mathrm{ad}(\mathfrak{so}(3))$-submodules
${\mathfrak{m}}={\mathfrak{m}}_{1}\oplus{\mathfrak{m}}_{2}\oplus{\mathfrak{m}}_{3}.$ (2.4)
The subalgebra $\mathfrak{so}(2)\oplus\mathfrak{so}(3)=\operatorname{{\rm Span}}\{e_{1},e_{8},e_{9},e_{10}\}\subset\mathfrak{so}(5)$ provides a reductive decomposition $\mathfrak{so}(5)=\mathfrak{so}(2)\oplus\mathfrak{so}(3)\oplus{\mathfrak{p}}$ (where ${\mathfrak{p}}={\mathfrak{m}}_{2}\oplus{\mathfrak{m}}_{3}$) for the symmetric space
$G_{2}(\mathbb{R}^{5})=\frac{{\rm SO}(5)}{{\rm SO}(2)\times{\rm SO}(3)},$ (2.5)
furthermore, ${\mathfrak{p}}\simeq\mathbb{C}^{3}$ with respect to the canonical almost complex structure $J=\begin{pmatrix}0&-I_{3}\\ I_{3}&0\end{pmatrix}$ and it turns out that ${\mathfrak{m}}_{3}=\operatorname{{\rm Span}}\{Je_{2},Je_{3},Je_{4}\}$. With respect to the basis (2.2), we define the alternating forms on ${\mathfrak{p}}$ given by
$\begin{array}[]{rl}\omega_{0}&=e^{25}+e^{36}+e^{47},\\[3.0pt] \mathop{\mathrm{Im}}\Omega_{0}&=e^{237}-e^{246}+e^{345}-e^{567},\\[3.0pt] \mathop{\mathrm{Re}}\Omega_{0}&=e^{234}-e^{267}+e^{357}-e^{456}.\end{array}$ (2.6)
Notice that $\omega_{0}\wedge\mathop{\mathrm{Re}}\Omega_{0}=0$ and $2\omega_{0}^{3}=3\mathop{\mathrm{Re}}\Omega_{0}\wedge\mathop{\mathrm{Im}}\Omega_{0}\neq 0$, i.e. $(\omega_{0},\mathop{\mathrm{Re}}\Omega_{0})$ is a ${\rm SU}(3)$-structure on ${\mathfrak{p}}$ (with induced orientation $-e^{234567}$).
Lemma 2.1.
The forms in (2.6) are $\mathrm{ad}(\mathfrak{so}(2)\oplus\mathfrak{so}(3))$-invariant, therefore, it induces a ${\rm SO}(5)$-invariant ${\rm SU}(3)$-structure $(\omega,\mathop{\mathrm{Re}}\Omega)$ on $G_{2}(\mathbb{R}^{5})$. Furthermore, we have that $d\omega=0,$ $d(\mathop{\mathrm{Re}}\Omega)=3e^{1}\wedge\mathop{\mathrm{Im}}\Omega$ and $d(\mathop{\mathrm{Im}}\Omega)=-3e^{1}\wedge\mathop{\mathrm{Re}}\Omega.$
Proof.
The Lie algebras $\mathfrak{sl}(\mathbb{C}^{3})$ and ${\mathfrak{s}}{\mathfrak{p}}(\mathbb{R}^{6})$ are described by:
$\mathfrak{sl}(\mathbb{C}^{3})=\left\{\begin{pmatrix}B&-C\\ C&B\end{pmatrix}\in\mathfrak{gl}({\mathfrak{p}})\Big{|}B,C\in\mathfrak{sl}(\mathbb{R}^{3})\right\},\quad{\mathfrak{s}}{\mathfrak{p}}(\mathbb{R}^{6})=\left\{\begin{pmatrix}B&C\\ D&-B^{T}\end{pmatrix}\in\mathfrak{gl}({\mathfrak{p}})\Big{|}C,D\in sym(\mathbb{R}^{3})\right\}.$
The Lie algebras $\mathfrak{sl}(\mathbb{C}^{3})$ and ${\mathfrak{s}}{\mathfrak{p}}(\mathbb{R}^{6})$ can be interpreted as stabilisers
$\mathfrak{stab}_{\mathfrak{gl}({\mathfrak{p}})}(\omega_{0})=\mathfrak{sl}(\mathbb{C}^{3})$
and
$\mathfrak{stab}_{\mathfrak{gl}({\mathfrak{p}})}(\mathop{\mathrm{Re}}\Omega_{0})=\mathfrak{stab}_{\mathfrak{gl}({\mathfrak{p}})}(\mathop{\mathrm{Im}}\Omega_{0})=\mathfrak{sp}(\mathbb{R}^{6}).$
Notice that the endomorphisms $\mathrm{ad}(e_{8})_{\mathfrak{p}}$, $\mathrm{ad}(e_{9})_{\mathfrak{p}}$, $\mathrm{ad}(e_{10})_{\mathfrak{p}}$ and $\mathrm{ad}(e_{1})_{\mathfrak{p}}=J$ belong to the interception $\mathfrak{sl}(\mathbb{C}^{3})\cap\mathfrak{sp}(\mathbb{R}^{6})$. Therefore, the forms in (2.6) are $\mathrm{ad}(\mathfrak{so}(2)\oplus\mathfrak{so}(3))$-invariant. Now, for any $p=x\cdot\Big{(}{\rm SO}(2)\times{\rm SO}(3)\Big{)}\in G_{5,2}$ define $(\omega,\real\Omega)$ by
$\omega_{p}={\rm L}_{x}^{\ast}\omega_{0}\quad\text{and}\quad\real\Omega_{p}={\rm L}_{x}^{\ast}\real\Omega_{0},$
where ${\rm L}_{x}$ is the left multiplication by $x\in{\rm SO}(5)$, hence $2\omega^{3}=3\mathop{\mathrm{Re}}\Omega\wedge\mathop{\mathrm{Im}}\Omega\neq 0$ Finally, from Table 1 follows that $d\omega_{0}=0,$ $d(\mathop{\mathrm{Re}}\Omega_{0})=3e^{1}\wedge\mathop{\mathrm{Im}}\Omega_{0}$ and $d(\mathop{\mathrm{Im}}\Omega_{0})=-3e^{1}\wedge\mathop{\mathrm{Re}}\Omega_{0}$. ∎
The following result is mentioned in [kerr1998new]*§4, for sake of completeness, we include the detail proof of it.
Lemma 2.2.
Any $\mathrm{Ad}({\rm SO}(3))$-invariant metric on $V^{5,2}$ is isometric to a diagonal one, i.e., up to isometries, it is enough to consider the class of invariant metrics determined by the $\mathrm{Ad}({\rm SO}(3))$-invariant inner product:
$g=y_{1}Q|_{{\mathfrak{m}}_{1}}+y_{2}Q|_{{\mathfrak{m}}_{2}}+y_{3}Q|_{{\mathfrak{m}}_{3}},\quad\text{where}\quad y_{i}>0\quad i=1,2,3.$ (2.7)
Proof.
In the decomposition (2.4) we have ${\mathfrak{m}}_{2}\simeq{\mathfrak{m}}_{3}$, hence, any ${\rm SO}(5)$-invariant metric on $V^{5,2}$ is determined by the $\mathrm{Ad}({\rm SO}(3))$-invariant inner product $\langle\cdot,\cdot\rangle=Q(g\cdot,\cdot)$, where $g$ is a positive symmetric matrix of the form
$g=\begin{pmatrix}y_{1}&0&0\\ 0&y_{2}I_{3}&xI_{3}\\ 0&xI_{3}&y_{3}I_{3}\end{pmatrix}\quad\text{for}\quad y_{1},y_{2},y_{3}>0,\quad x\in\mathbb{R}\quad\text{with}\quad y_{2}y_{3}>x^{2}.$
Denote by $N({\rm SO}(3))$ the normalizer of ${\rm SO}(3)$ in ${\rm SO}(5)$, since the quotient $\frac{N({\rm SO}(3))}{{\rm SO}(3)}$ is isomorphic to ${\rm SO}(2)$ the conjugation induces a 1-parameter family of diffeomorphism on ${\mathfrak{m}}$ given by $\phi(t)=\mathrm{Ad}(\exp(te_{1}))$. Thus, for $t\in\mathbb{R}$ the one-parameter family $g_{t}=\phi(t)^{T}g\phi(t)$ defines an isometric $\mathrm{Ad}({\rm SO}(3))$-inner product.
A straightforward computation shows that the parameter $x(t)$ in the metric $g_{t}$, is given by $x(t)=\frac{(y_{3}-y_{2})}{2}\sin(2t)+x\cos(2t)$. Hence, $g_{t}$ is diagonal if:
$y_{2}\neq y_{3}\quad\text{and}\quad t=\frac{1}{2}\arctan\left(\frac{2x}{y_{2}-y_{3}}\right)$
or
$y_{2}=y_{3}\quad\text{and}\quad t=\frac{(2k+1)\pi}{4}\quad\text{with}\quad k\in{\bf N},$
With respect to (2.7), we define an orthonormal basis of ${\mathfrak{m}}$:
$X_{1}:=\frac{1}{\sqrt{y_{1}}}e_{1},\quad X_{i}:=\frac{1}{\sqrt{y_{2}}}e_{i}\quad\text{and}\quad X_{i+3}:=\frac{1}{\sqrt{y_{3}}}e_{i+3}\quad\text{for}\quad i=1,2,3.$ (2.8)
Then, the Lie bracket in Table 1 becomes
$\begin{split}[X_{1},X_{i}]_{{\mathfrak{m}}}=-\sqrt{\frac{y_{3}}{y_{1}y_{2}}}X_{i+3},\quad[X_{1},X_{i+3}]_{{\mathfrak{m}}}=\sqrt{\frac{y_{2}}{y_{1}y_{3}}}X_{i},\quad[X_{i},X_{i+3}]_{{\mathfrak{m}}}=-\sqrt{\frac{y_{1}}{y_{2}y_{3}}}X_{1}.\end{split}$ (2.9)
In [reidegeld2010spaces], Reidegeld proved that the space of ${\rm SO}(5)$-invariant $\mathrm{G}_{2}$-structures on $V^{5,2}$ has dimension $5$. The next Lemma describes the invariant $\mathrm{G}_{2}$-structures on the Stiefel manifold up to $\mathrm{G}_{2}$-equivalence, i.e. for a $\mathrm{G}_{2}$-structure $\varphi$ written in the basis $\{e_{1},...,e_{7}\}$ and $\{\tilde{e}_{1},...,\tilde{e}_{7}\}$ of ${\mathfrak{p}}$, there is a change of basis $f\in\mathrm{G}_{2}$ from $\{e_{1},...,e_{7}\}$ to $\{\tilde{e}_{1},...,\tilde{e}_{7}\}$.
Lemma 2.3.
The $5$-dimensional family of ${\rm SO}(5)$-invariant $\mathrm{G}_{2}$-structures on $V^{5,2}$ are described, up to $\mathrm{G}_{2}$-equivalence, by $(a,b,x,y,z)\in\mathbb{R}^{5}$ satisfying
$a^{2}+by>0,\quad b^{2}+ax>0\quad z\neq 0\quad\text{and}\quad xy=ab.$ (2.10)
Explicitly, the $\mathrm{G}_{2}$-structure $\varphi$ can be written by
$\varphi=$ $-X^{1}\wedge(X^{25}+X^{36}+X^{47})+p_{1}X^{234}+p_{2}(-X^{267}+X^{357}-X^{456})$ (2.11)
$+p_{3}(X^{237}-X^{246}+X^{345})-p_{4}X^{567},$
where
$p_{1}=\frac{x\sqrt{a^{2}+by}}{b^{2}+ax},\quad p_{2}=\frac{a}{\sqrt{a^{2}+by}},\quad p_{3}=\frac{b}{\sqrt{b^{2}+ax}}\quad\text{and}\quad p_{4}=\frac{y\sqrt{b^{2}+ax}}{a^{2}+by}$ (2.12)
and
$y_{1}=\frac{z^{2}}{y_{2}y_{3}},\quad y_{2}=\frac{(b^{2}+ax)^{2/3}}{(a^{2}+by)^{1/3}}\quad\text{and}\quad y_{3}=\frac{(a^{2}+by)^{2/3}}{(b^{2}+ax)^{1/3}}.$ (2.13)
Proof.
From (2.6) is easy to check that any $\mathrm{Ad}({\rm SO}(3))$-invariant $3$-form on ${\mathfrak{m}}$ is given by
$\widetilde{\varphi}=$ $-\widetilde{z}e^{1}\wedge(e^{25}+e^{36}+e^{47})+\widetilde{x}e^{234}+\widetilde{a}(-e^{267}+e^{357}-e^{456})$
$+\widetilde{b}(e^{237}-e^{246}+e^{345})-\widetilde{y}e^{567}.$
The $3$-form $\widetilde{\varphi}$ defines a $\mathrm{G}_{2}$-structure, if the induced symmetric bilinear form
$B_{ij}=\Big{(}(e_{i}\lrcorner\varphi)\wedge(e_{j}\lrcorner\varphi)\wedge\varphi\Big{)}(e_{1},...,e_{7})$
is positive (or negative) defined. Explicitly, we obtain that for $k=2,3,4$
$B_{11}=6\widetilde{z}^{3},\quad B_{kk}=6\widetilde{z}(\widetilde{b}^{2}+\widetilde{a}\widetilde{x}),\quad B_{k+3k+3}=6\widetilde{z}(\widetilde{a}^{2}+\widetilde{b}\widetilde{x}),\quad B_{kk+3}=3\widetilde{z}(\widetilde{a}\widetilde{b}-\widetilde{x}\widetilde{y})$
and $B_{ij}=0$ otherwise.
Now, since $\phi(t)=\mathrm{Ad}(\exp(te_{1}))\in{\rm SU}(3)\subset\mathrm{G}_{2}$ [cf. Lemma 2.2], there exist $t_{0}\in\mathbb{R}$ such that
$\varphi=\phi(t_{0})^{*}\widetilde{\varphi}=$ $-ze^{1}\wedge(e^{25}+e^{36}+e^{47})+xe^{234}+a(-e^{267}+e^{357}-e^{456})$
$+b(e^{237}-e^{246}+e^{345})-ye^{567},$
satisfies (2.10). On the other hand, the induced metric is given by
$g=\frac{B}{6^{\frac{2}{9}}\mathop{\mathrm{missing}}{det}\nolimits(B)^{\frac{1}{9}}},$
hence, by comparing $g$ with (2.7), we obtain the condition in (2.13). Finally, the expression in (2.11) is obtained by writing $\varphi$ in the orthonormal basis (2.8). ∎
Definition 2.4.
A $\mathrm{G}_{2}$-structure $\varphi$ is called coclosed if $d\psi=0$. Among the class of coclosed class, $\varphi$ is said purely coclosed if $d\varphi\wedge\varphi=0$ and nearly parallel if $d\varphi=4\psi$.
From Lemma (2.3), the following result describe all coclosed invariant $\mathrm{G}_{2}$–metrics on $V^{5,2}$
Theorem 2.1.
The $\mathrm{G}_{2}$-structure (2.11) is coclosed for any $(a,b,x,y,z)\in\mathbb{R}^{5}$ satisfying (2.10). Moreover, $V^{5,2}$ does not carry a ${\rm SO}(5)$-invariant purely coclosed $\mathrm{G}_{2}$-structure and (2.11) is nearly parallel if
$x=a,\quad y=b,\quad a^{2}+b^{2}=\frac{27}{512}\quad\text{and}\quad z=-\frac{9}{32}.$ (2.14)
Proof.
From Lemma 2.3, we get the dual 4-form $\psi=\ast\varphi$
$\psi=$ $-(X^{2356}+X^{2457}+X^{3467})-p_{4}X^{1234}+p_{3}X^{1}\wedge(X^{267}-X^{357}+X^{456})$ (2.15)
$+p_{2}X^{1}\wedge(X^{237}-X^{246}+X^{345})-p_{1}X^{1567}.$
Using the Lie bracket (2.9) and the metric coefficients (2.7), we obtain the exterior derivative for the following $\mathrm{Ad}({\rm SO}(3))$-forms on ${\mathfrak{m}}$:
$dX^{1}=$ $\frac{z}{y_{2}y_{3}}\left(X^{25}+X^{36}+X^{47}\right)$ (2.16)
$dX^{234}=$ $-\frac{y_{2}}{z}\left(X^{1237}-X^{1246}+X^{1345}\right),\quad dX^{567}=\frac{y_{3}}{z}\left(X^{1267}-X^{1357}+X^{1456}\right)$
$d(X^{267}$ $-X^{357}+X^{456})=\frac{2y_{3}}{z}\left(X^{1237}-X^{1246}+X^{1345}\right)-\frac{3y_{2}}{z}X^{1567}$
$d(X^{237}$ $-X^{246}+X^{345})=\frac{3y_{3}}{z}X^{1234}-\frac{2y_{2}}{z}\left(X^{1267}-X^{1357}+X^{1456}\right).$
From (2.15) and (2.16), it is easy to check that $d\psi=0$. Now, from (2.11) we obtain
$d\varphi=$ $\frac{2z}{y_{2}y_{3}}\left(X^{2356}+X^{2457}+X^{3467}\right)-\frac{p_{1}y_{2}+2p_{2}y_{3}}{z}\left(X^{1237}-X^{1246}+X^{1345}\right)$ (2.17)
$+\frac{3p_{2}y_{2}}{z}X^{1567}+\frac{3p_{3}y_{3}}{z}X^{1234}-\frac{p_{4}y_{3}+2p_{3}y_{2}}{z}\left(X^{1267}-X^{1357}+X^{1456}\right).$
From a straightforward computation, we note that $\ast\left(d\varphi\wedge\varphi\right)=12+\frac{3}{y_{2}}+\frac{3}{y_{3}}\neq 0,$ in other words, $\varphi$ given in (2.11) is not purely coclosed. Finally, by imposing the condition $d\varphi=4\psi$, and using (2.15) and (2.17), we obtain (2.14). ∎
Remark 2.5.
The $S^{1}$-family given in (2.14) correspond to the general construction given [alexandrov2012]*(4.28).
2.2 Homogeneous Sasakian metrics on $V^{5,2}$
In this section, we study the Sasakian condition for ${\rm SO}(5)$-invariant metrics on $V^{5,2}$. Firstly, we characterise the class $(\eta,\xi,g,\Phi)$ of contact metrics structures, cf. Section 1. Secondly, in Theorem 2.2, we provide the constraints in the metric to get a Sasakian structures. And finally, we relate the invariant $\mathrm{G}_{2}$-structures inducing a Sasakian metric.
Lemma 2.6.
The globally defined $1$-form $\eta=X^{1}$ and $\xi=X_{1}$ define a ${\rm SO}(5)$-invariant almost contact structure on $V^{5,2}$. Furthermore, $(\eta,\xi,g,\Phi)$ defines a contact metric structure on $V^{5,2}$, if and only if
$g=4y_{2}y_{3}Q|_{{\mathfrak{m}}_{1}}+y_{2}Q|_{{\mathfrak{m}}_{2}}+y_{3}Q|_{{\mathfrak{m}}_{3}}\quad\text{and}\quad\Phi(X)=[e_{1},X].$ (2.18)
Proof.
First, note that $(V^{5,2},\eta)$ is a contact manifold, since it is clear that $\eta(\xi)=1$ and that $\eta\wedge d\eta^{3}\neq 0$ follows from (2.16). Now, we define an $\mathrm{Ad}({\rm SO}(3))$-invariant contact metric structure $(\eta,\xi,g,\Phi)$, i.e., we define an $\mathrm{Ad}({\rm SO}(3)$-invariant endomorphism $\Phi$ on ${\mathfrak{m}}$ such that $\Phi^{2}=-I+\eta\otimes\xi$ and $d\eta(X,Y)=2g(X,\Phi(Y))$. From (2.9) we have
$d\eta=\sqrt{\frac{y_{1}}{y_{2}y_{3}}}\left(X^{25}+X^{36}+X^{47}\right),$
Hence, the condition $d\eta(X_{i},X_{i+3})=2g(X_{i},\Phi(X_{i+3}))$ implies that $\Phi$ is given by
$\Phi X_{i}=-\frac{1}{2}\sqrt{\frac{y_{1}}{y_{2}y_{3}}}X_{i+3},\quad\Phi X_{i+3}=\frac{1}{2}\sqrt{\frac{y_{1}}{y_{2}y_{3}}}X_{i},\quad\text{and}\quad\Phi\xi=0.$ (2.19)
Moreover, using the above equalities and the condition $\Phi^{2}=-I+\eta\otimes\xi$, we obtain that $-\frac{1}{4}\frac{y_{1}}{y_{2}y_{3}}=-1$, or equivalently $y_{1}=4y_{2}y_{3}$. ∎
The Nijenhuis tensor induced by $\Phi$ is defined by
$N_{\Phi}(X,Y)=[\Phi X,\Phi Y]+\Phi^{2}[X,Y]-\Phi[\Phi X,Y]-\Phi[X,\Phi Y].$ (2.20)
From [Boyer2008]*Theorem 6.5.9, we know that an almost contact structure $(\xi,\eta,\Phi)$ is normal if the Nijenhuis tensor satisfies
$N_{\Phi}=-d\eta\otimes\xi.$ (2.21)
Following [Boyer2008]*Definition 6.5.13, we define:
Definition 2.7.
A normal contact metric structure $\mathcal{S}=(\eta,\xi,g,\Phi)$ on $M^{2n+1}$ is called Sasakian and $(M^{2n+1},\mathcal{S})$ is called a Sasakian manifold.
Theorem 2.2.
The ${\rm SO}(5)$-invariant contact metric structure $(\eta,\xi,g,\Phi)$ given in Lemma 2.6 is a Sasakian structure, if and only if, the metric (2.18) satisfies
$y_{1}=4y_{2}^{2}\quad\text{and}\quad y_{2}=y_{3}.$ (2.22)
Proof.
We show that $(\eta,\xi,g,\Phi)$ is a normal, contact metric structure, i.e., Sasakian. Using the bracket (2.9) and the definition of $\Phi$ in (2.18), we verify that the Nijenhuis tensor (2.20) satisfies (2.21). From Lemma 2.6, we recall that the metric $g$ is contact for $y_{1}=4y_{2}y_{3}$. First, note that $N_{\Phi}(\xi,Y)=\Phi^{2}[\xi,Y]-\Phi[\xi,\Phi Y]$ and now, for $i=1,2,3$ we consider
$\Phi^{2}[\xi,X_{i}]-\Phi[\xi,\Phi X_{i}]$ $=-[\xi,X_{i}]-\Phi[\xi,-X_{i+3}]=-[\xi,X_{i}]+\sqrt{\frac{y_{2}}{y_{1}y_{3}}}\Phi X_{i}$
$=\sqrt{\frac{y_{3}}{y_{1}y_{2}}}X_{i+3}-\sqrt{\frac{y_{2}}{y_{1}y_{3}}}X_{i+3}$
$=\left(\frac{1}{2y_{2}}-\frac{1}{2y_{3}}\right)X_{i+3}$
Since $i_{\xi}d\eta=0$, we conclude that $y_{2}=y_{3}$. Analogously
$\Phi^{2}[\xi,X_{i+3}]-\Phi[\xi,\Phi X_{i+3}]$ $=-[\xi,X_{i+3}]-\Phi[\xi,X_{i}]=\sqrt{\frac{y_{2}}{y_{1}y_{3}}}X_{i}-\sqrt{\frac{y_{3}}{y_{1}y_{2}}}\Phi X_{i+3}$
$=\frac{1}{\sqrt{y_{1}}}X_{i}-\frac{1}{\sqrt{y_{1}}}X_{i}=0.$
For the missing cases, for $X,Y\in{\mathfrak{m}}_{2}\oplus{\mathfrak{m}}_{3}$ the brackets $[X,Y]$, $[\Phi Y,Y]$ and $[X,\Phi Y]$ are generated by $\xi$ (cf. (2.9)), hence $\Phi[X,Y]=\Phi[\Phi X,Y]=\Phi[X,\Phi Y]=0$ and thus $N_{\Phi}(X,Y)=[\Phi X,\Phi Y]$. So, (2.21) holds by noting that
$N_{\Phi}(X_{i},X_{i+3})=-2X_{1}\quad\text{for}\quad i=2,3,4\quad\text{and}\quad N_{\Phi}(X_{i},X_{j})=0\quad\text{otherwise}.$
According to the notation in Theorem 2.2, the Sasakian metric (2.7) is induced by the following family of $\mathrm{G}_{2}$-structures.
Corollary 2.8.
Consider the Sasakian metric on $V^{5,2}$ as a metric induced by (2.11) and the parameter vector $(z,a,b,x,y)\in\mathbb{R}^{5}$ satisfying the condition (2.10). Then, the $\mathrm{G}_{2}$-structure (2.11) induces a Sasakian metric, if and only if,
$x=a,\quad y=b\quad\text{and}\quad z^{2}=4(a^{2}+b^{2})^{\frac{4}{3}}.$ (2.23)
In this case, the $2$-parameter family of coclosed $\mathrm{G}_{2}$-structures is
$\varphi=$ $-X^{1}\wedge(X^{25}+X^{36}+X^{47})+\left(\frac{2}{z}\right)^{\frac{3}{4}}\Big{\{}a\Big{(}X^{234}-X^{267}+X^{357}-X^{456}\Big{)}$
$+b\Big{(}X^{237}-X^{246}+X^{345}-X^{567}\Big{)}\Big{\}},$
$\psi=$ $-\Big{(}X^{2356}+X^{2457}+X^{3467}\Big{)}-\left(\frac{2}{z}\right)^{\frac{3}{4}}\Big{\{}bX^{1}\wedge\Big{(}X^{234}-X^{267}+X^{357}-X^{456}\Big{)}$
$-aX^{1}\wedge\Big{(}X^{237}-X^{246}+X^{345}-X^{567}\Big{)}\Big{\}}.$
Proof.
To proof item $(1)$, we compare (2.13) and (2.22), thus
$z^{2}=4(a^{2}+by)^{4/3}\quad\text{and}\quad a^{2}+by=b^{2}+ax>0,$
since $a^{2}b^{2}=abxy$, we have $a^{2}(ax+b^{2})=(b^{2}+ax)ax$ and $by(a^{2}+by)=b^{2}(by+a^{2})$. Thus, the result follows.
Remark 2.9.
On the $\mathrm{ad}(\mathfrak{so}(3))$-invariant reductive decomposition ${\mathfrak{m}}={\mathfrak{m}}_{1}\oplus{\mathfrak{m}}_{2}\oplus{\mathfrak{m}}_{3}$ (2.3) the tensors given by
$\xi=\frac{4}{3}e_{1},\quad\Phi=\begin{pmatrix}0&&\\ &&I_{3}\\ &-I_{3}&\end{pmatrix}\quad\text{and}\quad g=\frac{3}{8}\begin{pmatrix}\frac{3}{8}&0\\ 0&I_{6}\end{pmatrix},$
define the ${\rm SO}(5)$-invariant Sasakian-Einstein structure on $V^{5,2}$, associated with nearly parallel $\mathrm{G}_{2}$-structure of type II (cf.[friedrich1997nearly]*Table 2).
3 $\mathrm{G}_{2}$ and self dual contact instantons on $V^{5,2}$
According to Corollary 2.8, we can consider solutions of (1.2) for those $\mathrm{G}_{2}$-structures with induced Sasakian metric. Additionally, if we endow a complex vector bundle $E\to V^{5,2}$ with a holomorphic Sasakian structure [biswas2010vector], one obtain the notion of Chern connection, as those connections simultaneously compatible with the holomorphic structure and some Hermitian bundle metric. The Hermitian Yang-Mills (HYM) condition is naturally extended to a Sasakian version, namely, let $\omega\colon=d\eta\in\Omega^{1,1}(M)$ be the transversal Kähler form, we say that a connection $A\in\mathcal{A}(E)$ is transverse Hermitian Yang-Mills connection (tHYM) if
$\hat{F}_{A}\colon=\langle F_{A},\omega\rangle=0,\quad\text{and}\quad F^{0,2}_{A}=0.$ (3.1)
Here, the type $(p,q)$ of a $k$-form on $M$ is induced by the condition $\Phi|_{H}^{2}=-1$, where $H$ is the Kähler distribution $H=\ker\eta\subset TM=H\oplus\xi$. Additionally, the splitting on the tangent bundle induced by the contact structure yields the splitting $\Omega^{k}(M)=\Omega^{k}(H^{\ast})\oplus\eta\wedge\Omega^{k-1}(H^{\ast})$. Hence, after complexification of the tangent bundle, one can naturally extend this bi-degree $(p,q)$ to the bundle of $E$-valued $k$-forms. We know that $2$-forms on $M$, and consequently $\Omega^{2}({\mathfrak{g}}_{E})$, decomposes by (cf. [portilla2023, Section. 2.1])
$\Omega^{2}(M)=\underbrace{\Omega^{2}_{1}\oplus\Omega^{2}_{6}\oplus\Omega^{2}_{8}}_{\text{Transversal $2$-forms}}\oplus\Omega^{2}_{V}$ (3.2)
Furthermore, a connection $A\in\mathcal{A}(E)$ is SDCI if its curvature belong to $\Omega^{2}_{8}({\mathfrak{g}}_{E})$, or equivalently, if the curvature $F_{A}\in\Omega^{2}({\mathfrak{g}}_{E})$ is of type $(1,1)$ orthogonal to $\omega$ (cf. [portilla2023, Proposition 2.5]).
Lemma 3.1.
With respect to the Sasakian structure of $V^{5,2}$ given in Theorem 2.2 and the $\mathrm{G}_{2}$-structure in (2.11) we have that: If $A$ is a SDCI, then it is a $\mathrm{G}_{2}$-instanton.
Proof.
We know that in local coordinates, a basis for self dual contact $2$–forms is given by (cf. [portilla2023, Section 2])
$\begin{array}[]{lccl}w_{1}=X^{23}+X^{56},&w_{2}=X^{26}+X^{35},&w_{3}=X^{24}+X^{57},&w_{4}=X^{27}+X^{45},\\[4.0pt] w_{5}=X^{34}+X^{67},&w_{6}=X^{37}+X^{46},&w_{7}=X^{25}-X^{47},&w_{8}=X^{36}-X^{47}.\end{array}$ (3.3)
For $\mathop{\mathrm{Re}}\Omega_{0}$ and $\mathop{\mathrm{Im}}\Omega_{0}$ given in Lemma 2.1, one verifies that $w_{i}\wedge\real\Omega_{0}=w_{i}\wedge\mathop{\mathrm{Im}}\Omega_{0}=0$ for each $i=1,\cdots,8$ in (3.3). Hence, if $A$ is a SDCI, then its curvature $F_{A}$ is locally generated by $w_{i}$’s in (3.3). Thus, for the $\mathrm{G}_{2}$-structure (cf. Corollary 2.8):
$\varphi=-\eta\wedge d\eta+\left(\frac{2}{z}\right)^{\frac{3}{4}}a\mathop{\mathrm{Re}}(\Omega_{0})+\left(\frac{2}{z}\right)^{\frac{3}{4}}b\mathop{\mathrm{Im}}(\Omega_{0}),$
we have
$\ast(\varphi\wedge F_{A})$ $=-\ast(\eta\wedge d\eta\wedge F_{A})+\left(\frac{2}{z}\right)^{\frac{3}{4}}a\ast(\mathop{\mathrm{Re}}(\Omega_{0})\wedge F_{A})+\left(\frac{2}{z}\right)^{\frac{3}{4}}b\ast(\mathop{\mathrm{Im}}(\Omega_{0})\wedge F_{A})$
$=-\ast(\eta\wedge d\eta\wedge F_{A})=-F_{A}.$
Remark 3.2.
We can generalise Lemma 3.1 in the following sense. Observe that from the decomposition in (3.2), we have the ${\rm SU}(3)$-invariant decomposition
$\Omega^{2}(H)=\Omega^{2}_{6}\oplus\Omega^{2}_{1}\oplus\Omega_{8}^{2},$ (3.4)
where $\Omega^{2}_{8}\simeq\mathfrak{su}(3)$ and $\Omega^{2}_{1}=\Omega^{0}(M)\otimes\omega$. On the other hand, we know that the space of $2$-forms on $H^{\mathbb{C}}=H\otimes_{\mathbb{R}}\mathbb{C}$ decomposes into $\operatorname{Gl}(3,\mathbb{C})$-invariant submodules
$\Omega^{2}(H^{\mathbb{C}})=\Omega^{2,0}(H^{\mathbb{C}})\oplus\Omega^{0,2}(H^{\mathbb{C}})\oplus\Omega^{1,1}(H^{\mathbb{C}}).$
By Schur’s Lemma, we have (see [falcitelli1994]*Equation (1.2)
$\Omega^{2,0}\oplus\Omega^{0,2}\simeq\Omega^{2}_{6}\otimes_{\mathbb{R}}\mathbb{C}\quad\text{and}\quad\Omega^{1,1}\simeq(\Omega^{2}_{1}\oplus\Omega^{2}_{8})\otimes_{\mathbb{R}}\mathbb{C}.$ (3.5)
On the other hand, we have the $\mathrm{G}_{2}$-invariant decomposition $\Omega^{2}(TM)=\Omega^{2}_{7}\oplus\Omega^{2}_{14},$ where $\Omega^{2}_{14}\simeq{\mathfrak{g}}_{2}$. Since ${\rm SU}(3)\subset\mathrm{G}_{2}$ we have that $\Omega_{8}^{2}\subset\Omega_{14}^{2}$ and $\Omega^{2}_{1}\subset\Omega^{2}_{7}$. It is easy to see that $\Omega^{2}_{V}$ has elements in both $\Omega_{7}^{2}$ and $\Omega_{14}^{2}$, hence the same is true for $\Omega^{2}_{6}$.
Lemma 3.3.
Let $E\to V^{5,2}$ be a Sasakian holomorphic vector bundle [Boyer2008, Section 3], with the Sasakian structure $(\eta,\xi,g,\Phi)$ given in Theorem 2.2 and $A\in\mathcal{A}(E)$ the Chern connection. If $A\in\mathcal{A}(E)$ is a $\mathrm{G}_{2}$-instanton then $A$ satisfies the equation (1.2), i.e. it is a $\lambda$-contact instanton .
Proof.
Consider $\varphi$ given in Corollary 2.8 and set $dz^{k}=X^{k}+\sqrt{-1}X^{k+3}$ for $k=2,3,4$. From Lemma 2.1 and (2.6), we know that $\Omega_{0}=dz^{2}\wedge dz^{3}\wedge dz^{4}\in\Lambda^{3,0}H^{\ast}_{\mathbb{C}}$, where $H:=\ker(\eta)$. Now, if $A\in\mathcal{A}(E)$ is the Chern connection, then $F_{A}$ is of $(1,1)$-type [biswas2010vector, p.552]. By taking into consideration the bi-degree of $\Omega_{0}$, we conclude that $F_{A}\wedge\Omega_{0}=0$ and $F_{A}\wedge\overline{\Omega_{0}}=0$, hence both $F_{A}\wedge\mathop{\mathrm{Im}}{\Omega_{0}}$ and $F_{A}\wedge\mathop{\mathrm{Re}}{\Omega_{0}}$ vanishes, then
$F_{A}\wedge\varphi$ $=F_{A}\wedge\left(\eta\wedge d\eta+\left(\frac{2}{z}\right)^{\frac{3}{4}}a\mathop{\mathrm{Re}}\Omega_{0}+\left(\frac{2}{z}\right)^{\frac{3}{4}}b\mathop{\mathrm{Im}}\Omega_{0}\right)$
$=F_{A}\wedge\eta\wedge d\eta.$
Consequently, solutions to the $\mathrm{G}_{2}$-instanton equation are solutions to the contact instanton equation. ∎
Consider the Kähler manifold $G_{2}(\mathbb{R}^{5})\cong V^{5,2}/S^{1}$, we fix a holomorphic bundle $(\mathcal{E},\bar{\partial})$ on $G_{5,2}$, the pullback $\mathcal{E}^{\ast}:=\pi^{\ast}\mathcal{E}\to V^{5,2}$ carries a structure of Sasakian holomorphic bundle
[…]
where $\pi\colon V^{5,2}\to G_{2}(\mathbb{R}^{5})$ is a Sasakian circle fibration endowed with the $\mathrm{G}_{2}$-structure (2.11). Finally, in [portilla2023, Lemma 2.4] was proved that SDCI are precisely connections whose curvature is of type $(1,1)$ and orthogonal to $\omega$, (cf. the splitting in (3.2) and Remark 3.2). Thus, from Lemmas 3.1 and 3.3, we obtain the following result:
Corollary 3.4.
The vector bundle $\mathcal{E}^{\ast}$ is a holomorphic Sasakian bundle, and a Chern connection $A$ on $\mathcal{E}$ is HYM if, and only if, $\pi^{\ast}A$ is a $\mathrm{G}_{2}$-instanton on $\mathcal{E}^{*}$.
We recall that in Lemma 3.1, Lemma 3.3 and Corollary 3.4 the key argument follow just in presence of a SU$(n)$-structure. These have analogous results in [portilla2023, Proposition 3.11] and [calvo2020gauge, Lemma 21], the difference in our case is that we do not assume $M$ is a contact Calabi-Yau but only endowed with a transversal ${\rm SU}(3)$-structure.
4 Homogeneous principal bundles on $V^{5,2}$
First, we recall some notation and results about homogeneous principal bundle, following [cap2009]*Section 1.4. Let $H\subset G$ be a closed subgroup of the Lie group $G$, we denote by $M=G/H$ the homogeneous space, $L\colon G\times M\to M$ the canonical left action and $p\colon G\to M$ an $H$–principal bundle given by the canonical projection.
Definition 4.1.
A homogeneous principal bundle over $M$ is a locally trivial principal bundle $\pi\colon P\to M$ together with a left $G$–action $\widetilde{L}\colon G\times P\to P$, which lifts the action $L$ on $M$, i.e., $\pi(\widetilde{L}({\textsl{g}},y))=L({\textsl{g}},\pi(y))$, and for each ${\textsl{g}}\in G$, the bundle map $\widetilde{L}_{\textsl{g}}\colon P\to P$ is a homomorphism of principal bundles.
Homogeneous principal bundles are characterised in the following way:
Lemma 4.2.
[cap2009, Lemma 1.4.5] Let $G$ and $K$ be Lie groups and $H\subset G$ a closed Lie subgroup. Let $P\to G/H$ a homogeneous $K$-principal bundle, then there exist a smooth homomorphism $\phi\colon H\to K$ such that $P\cong G\times_{\phi}K$. The action of $H$ on $K$ is given by $h\cdot k=\phi(h)\cdot k$ for each $h\in H$ and $k\in K$. The bundles corresponding to two homomorphisms $\phi,\widetilde{\phi}\colon H\to K$ are isomorphic if and only if there exist $k\in K$ such that $k\cdot\phi=\widetilde{\phi}\cdot k$
For a homogeneous $K$-principal bundle $P\to G/H$, a connection $A\in\Omega^{1}(P,{\mathfrak{k}})$ is called invariant, if $\widetilde{L}_{\textsl{g}}^{*}A=A$ for each ${\textsl{g}}\in G$.
Lemma 4.3.
[cap2009, Theorem 1.4.5] Let $P=G\times_{\phi}K\to G/H$ be the homogeneous $K$-principal bundle induced by the homomorphism $\phi:H\to K$. Then, invariant principal connections on $P$ are in bijective correspondence with linear maps $\alpha\colon{\mathfrak{g}}\to{\mathfrak{k}}$ such that
  • (i)
    $\alpha|_{{\mathfrak{h}}}=\phi_{\ast}\colon{\mathfrak{h}}\to{\mathfrak{k}}$, the derivative of $\phi$.
  • (ii)
    $\alpha\circ\mathrm{Ad}(h)=\mathrm{Ad}(\phi(h))\circ\alpha$ for all $h\in H$.
4.1 Homogeneous ${\rm SO}(3)$-principal bundles
Consider $H=K={\rm SO}(3)$ and $G={\rm SO}(5)$. Since ${\rm SO}(3)$ is simple, up to conjugation there are two smooth homomorphisms $\phi_{i}\colon{\rm SO}(3)\to{\rm SO}(3),$ $i=0,1$. Namely, $\phi_{0}$ the trivial homomorphism and $\phi_{1}=1_{{\rm SO}(3)}$ the identity, according to Lemma 4.2, there exist precisely two [^3] Up to isomorphism of homogeneous bundles homogeneous ${\rm SO}(3)$-principal bundles $\pi_{i}\colon P_{i}=G\times_{\phi_{i}}{\rm SO}(3)\to V^{5,2}$, $i=0,1$.
Lemma 4.4.
There is a unique invariant ${\rm SO}(3)$–connection on $P_{0}=V^{5,2}\times{\rm SO}(3)$, given by
$\alpha=e^{1}\otimes(b_{8}e_{8}+b_{9}e_{9}+b_{10}e_{10}).$ (4.1)
Furthermore, the curvature $F_{\alpha}$ of the connection $\alpha$ belongs to $\Omega^{2}_{1}(\mathfrak{so}(3))$ cf. (3.4) and Remark 3.2.
Proof.
Consider an arbitrary linear map $\alpha:\mathfrak{so}(5)\rightarrow\mathfrak{so}(3)$, in the basis (2.2) we write
$\alpha=a_{i}e^{i}\otimes e_{8}+b_{j}e^{j}\otimes e_{9}+c_{k}e^{k}\otimes e_{10}.$
From Lemma 4.3, $\alpha$ is an invariant connection, if and only if, it satisfies $\alpha|_{\mathfrak{so}(3)}=0$ and $\alpha\circ\mathrm{Ad}({\textsl{g}})=\alpha$, for all ${\textsl{g}}\in{\rm SO}(3)$. The condition $\alpha|_{\mathfrak{so}(3)}=0$ implies that $\alpha$ is of the form $\alpha=a_{i}e^{i}\otimes e_{8}+b_{j}e^{j}\otimes e_{9}+c_{k}e^{k}\otimes e_{10}$ for $i,j,k\in\{1,\dots,7\}$. Now, we apply the second condition to ${\textsl{g}}_{k}=\begin{pmatrix}I_{2\times 2}&\\ 0&h_{k}\\ \end{pmatrix}\in{\rm SO}(5),$ we define some values of $h_{k}$,
$h_{0}=\begin{pmatrix}0&1&0\\ 1&0&0\\ 0&0&-1\end{pmatrix},h_{1}=\begin{pmatrix}0&1&0\\ -1&0&0\\ 0&0&1\end{pmatrix},h_{2}=\begin{pmatrix}0&0&1\\ 0&-1&0\\ 1&0&0\end{pmatrix},\quad\text{and}\quad h_{3}=\begin{pmatrix}0&0&1\\ 0&1&0\\ -1&0&0\end{pmatrix}$ (4.2)
First, considering $h_{0}$ and $h_{1}$, we obtain that $\alpha(e_{i})=a_{i}e_{8}+b_{i}e_{9}+c_{i}e_{10}$ and, for $i\neq 3$ and $i\neq 6$
$\alpha(\mathrm{Ad}({\textsl{g}}_{0})(e_{i}))=\alpha({\textsl{g}}_{0}e_{i}{\textsl{g}}_{0}^{-1})$ $=a_{i}e_{8}+b_{i}e_{10}+c_{i}e_{9}$
$\alpha(\mathrm{Ad}({\textsl{g}}_{1})(e_{i}))=\alpha({\textsl{g}}_{1}e_{i}{\textsl{g}}_{1}^{-1})$ $=-a_{i}e_{8}-b_{i}e_{10}-c_{i}e_{9}$
follows that $a_{i}=b_{i}=c_{i}=0$ for $i\in\{2,4,5,7\}$. Now, by considering $h_{2}$ and $h_{3},$ we obtain for $i=3,6$
$\begin{array}[]{cl}\alpha(\mathrm{Ad}({\textsl{g}}_{0})(e_{i}))=&-\alpha(e_{i})=a_{i}e_{8}+b_{i}e_{10}+c_{i}e_{9}\\[5.0pt] \alpha(\mathrm{Ad}({\textsl{g}}_{1})(e_{i}))=&\alpha(e_{i})=-a_{i}e_{8}-b_{i}e_{10}-c_{i}e_{9}.\end{array}$
Consequently, an invariant connection $\alpha$ is of the form
$\alpha=e^{1}\otimes(b_{8}e_{8}+b_{9}e_{9}+b_{10}e_{10}).$ (4.3)
The curvature of $\alpha$ is $F_{\alpha}=y_{2}^{-1}(X^{25}+X^{36}+X^{47})\otimes(b_{8}e_{8}+b_{9}e_{9}+b_{10}e_{10})$. The second affirmation in the Lemma follows from (2.6) since
$F_{\alpha}=2\omega\otimes(b_{8}e_{8}+b_{9}e_{9}+b_{10}e_{10}).$ (4.4)
Remark 4.5 (Homogeneous ${\rm U}(1)$-connections).
Similar to Lemma 4.4, since ${\rm SO}(3)$ is simple, all non-trivial homomorphism $\phi\colon{\rm SO}(3)\rightarrow K$ is either trivial or injective for all Lie group $K$. In particular for $K={\rm U}(1)$, from Lemma 4.2, there exists a unique [^4] Up to isomorphism of principal bundles homogeneous ${\rm U}(1)$-bundle $P$ corresponding to $\phi$, namely, the trivial bundle $P=V^{5,2}\times{\rm U}(1)$. Any invariant connection on $P$ is given by a linear map $\beta:\mathfrak{so}(5)\rightarrow{\mathfrak{u}}(1)$ satisfying
$(i)\quad\beta|_{\mathfrak{so}(3)}=0=\phi_{\ast}\colon\mathfrak{so}(3)\to\mathbb{R}\quad\text{and}\quad(ii)\quad\beta\circ\mathrm{Ad}(a)=\beta,\text{ for all }a\in{\rm SO}(3).$
By considering $e_{1}$ the generator of the Abelian Lie algebra ${\mathfrak{u}}(1)=\mathbb{R}$, the unique invariant connection is given by
$\alpha=a_{i}e^{i}\otimes e_{1}\quad\text{for}\quad i=1,\dots,10.$
From item $(i)$, since $\alpha|_{\mathfrak{so}(3)}=0$, $\alpha$ has to be of the form $\alpha=a_{i}e^{i}\otimes e_{1}$, $i=1,\cdots,7.$ On the other hand, considering item $(ii)$ for ${\textsl{g}}_{0},{\textsl{g}}_{1}$ given in (4.2), we use the same argument used in the proof of Lemma 4.4 to obtain $\alpha(e_{i})=a_{i}e_{1}$ and
$\alpha(Ad({\textsl{g}}_{0})(e_{i}))=\alpha({\textsl{g}}_{0}e_{i}{\textsl{g}}_{0}^{-1}))$ $=-a_{i}e_{1}$
$\alpha(Ad({\textsl{g}}_{1})(e_{i}))=\alpha({\textsl{g}}_{1}e_{i}{\textsl{g}}_{1}^{-1}))$ $=a_{i}e_{1}.$
for $i=2,\cdots,7$. Therefore, $\alpha$ is characterised by
$\beta=a_{1}e^{1}\otimes e_{1}$ (4.5)
and its curvature $2$-form is
$F_{\beta}=2(X^{25}+X^{36}+X^{47})\otimes e_{1}.$ (4.6)
Note that $F_{\beta}\in\Omega^{2}_{1}=\omega\otimes\Omega^{0}(M)$ (4.13).
Now, we consider the ${\rm SO}(3)$-principal bundle $P_{1}={\rm SO}(5)\times_{\phi_{1}}{\rm SO}(3)$ induced by the homomorphism identity $\phi_{1}\colon{\rm SO}(3)\to{\rm SO}(3)$. The projection of $P_{1}$ is given by
$\pi_{1}\colon[(u,s)]\in P_{1}\mapsto u\cdot{\rm SO}(3)\in V^{5,2},$ (4.7)
and the class is given by $[(u,s)]=\{(u\cdot r,r^{-1}s)\;|\;r\in{\rm SO}(3)\}.$ There exist a left ${\rm SO}(5)$-action on $P_{1}$ such that for each $v\in{\rm SO}(5)$ the bundle map $L_{v}\colon P_{1}\to P_{1}$ is $\pi_{1}$-equivariant i.e.
$\pi_{1}(L_{v}[(u,s)])=\pi_{1}([(vu,s)])=vu\cdot{\rm SO}(3)=L_{v}(\pi_{1}[(u,s)]).$
According to (2.2), we fix the basis $\{e_{8},e_{9},e_{10}\}$ for $\mathfrak{so}(3)\subset\mathfrak{so}(5)$ such that $[e_{i},e_{j}]=\varepsilon_{ijk}e_{k}$, where $\varepsilon_{ijk}$ is the sign of the permutation $(i\;j\;k)$.
Lemma 4.6.
There is a unique invariant ${\rm SO}(3)$–connection on $P_{1}=V^{5,2}\times_{\phi_{1}}{\rm SO}(3)$ given by
$\alpha=e^{8}\otimes e_{8}+e^{9}\otimes e_{9}+e^{10}\otimes e_{10}.$ (4.8)
Proof.
Consider $\alpha:\mathfrak{so}(5)\rightarrow\mathfrak{so}(3)$ in the basis (2.2) given by $\alpha=a_{i}e^{i}\otimes e_{8}+b_{j}e^{j}\otimes e_{9}+c_{k}e^{k}\otimes e_{10}$. From Lemma 4.3 (i), we have that $\alpha=\alpha_{0}+e^{8}\otimes e_{8}+e^{9}\otimes e_{9}+e^{10}\otimes e_{10},$ where $\alpha_{0}=a_{i}e^{i}\otimes e_{8}+b_{j}e^{j}\otimes e_{9}+c_{k}e^{k}\otimes e_{10}$, $i,j,k\in\{1,\dots,7\}$, we are going to check that $\alpha_{0}=0$. From Lemma 4.3 (ii), for any ${\textsl{g}}\in{\rm SO}(3)$ and $X\in\mathfrak{so}(5)$, we have ${\textsl{g}}\alpha(X){\textsl{g}}^{-1}=\alpha({\textsl{g}}X{\textsl{g}}^{-1})$. Taking $X=e_{i}$ for each $i\in\{1,\dots,7\}$, we denote
$\alpha(\mathrm{Ad}({\textsl{g}})(e_{i}))=\alpha({\textsl{g}}e_{i}{\textsl{g}}^{-1})$ (4.9)
and
$\mathrm{Ad}({\textsl{g}})(\alpha(e_{i}))=a_{i}{\textsl{g}}e_{8}{\textsl{g}}^{-1}+b_{i}{\textsl{g}}e_{9}{\textsl{g}}^{-1}+c_{i}{\textsl{g}}e_{10}{\textsl{g}}^{-1}.$ (4.10)
Now, consider ${\textsl{g}}_{k}$ and $h_{k}$ as in (4.2), equations (4.9) and (4.10) become
$\alpha(\mathrm{Ad}({\textsl{g}}_{0})(e_{2}))=\alpha({\textsl{g}}_{0}e_{2}{\textsl{g}}_{0}^{-1})=\alpha(e_{3})=a_{3}e_{8}+b_{3}e_{9}+c_{3}e_{10},$
and
$\mathrm{Ad}({\textsl{g}}_{0})(\alpha(e_{2}))=-a_{2}e_{8}+b_{2}e_{10}-c_{2}e_{9}.$
These imply $-a_{2}=a_{3},-c_{2}=b_{3},c_{3}=b_{2}$. Analogously, for $h_{1}$, we have
$\alpha(\mathrm{Ad}({\textsl{g}}_{1})(e_{2}))=\alpha(e_{3})=-a_{3}e_{8}-b_{3}e_{9}-c_{3}e_{10}$
and
$\mathrm{Ad}({\textsl{g}}_{1})(\alpha(e_{2}))=a_{2}e_{8}+b_{2}e_{10}-c_{2}e_{9},$
consequently, $-a_{2}=a_{3},c_{2}=b_{3},-b_{2}=c_{3}$, hence $c_{2}=b_{3}=c_{3}=b_{2}=0$. Similarly:
  • From $\mathrm{Ad}({\textsl{g}}_{0})(\alpha(e_{4}))=\alpha(\mathrm{Ad}({\textsl{g}}_{0})(e_{4}))$, we obtain $c_{4}=b_{4}=0$,
  • From $\mathrm{Ad}({\textsl{g}}_{1})(\alpha(e_{7}))=\alpha(\mathrm{Ad}({\textsl{g}}_{1})(e_{7}))$, we obtain $c_{7}=b_{7}=0$.
  • From $\mathrm{Ad}({\textsl{g}}_{0})(\alpha(e_{5}))=\alpha(\mathrm{Ad}({\textsl{g}}_{0})(e_{5}))$ and $\mathrm{Ad}({\textsl{g}}_{1})(\alpha(e_{5}))=\alpha(\mathrm{Ad}({\textsl{g}}_{1})(e_{5}))$, we obtain $c_{6}=b_{5}=c_{5}=b_{6}=0$.
Finally, applying the same argument for $h_{2}$ and $h_{3},$ we conclude that $a_{i}=0$ for $i=1,\cdots,7$, consequently $\alpha_{0}=0$, then $\alpha=e^{8}\otimes e_{8}+e^{9}\otimes e_{9}+e^{10}\otimes e_{10}$ and (4.8) defines a unique ${\rm SO}(5)$-invariant connection $\alpha\in\Omega^{1}(P,\mathfrak{so}(3))$. ∎
In the basis (2.8), the curvature of (4.8) is given by
$F_{\alpha}=\frac{1}{y_{2}}\left((X^{23}+X^{56})\otimes e_{8}+(X^{24}+X^{57})\otimes e_{9}+(X^{34}+X^{67})\otimes e_{10}\right).$ (4.11)
Notice that $F_{\alpha}:{\mathfrak{m}}\times{\mathfrak{m}}\rightarrow\mathfrak{so}(3)$ is a ${\rm SO}(3)$-equivariant map, i.e., $F_{\alpha}\circ\mathrm{Ad}({\textsl{g}})=\mathrm{Ad}({\textsl{g}})\circ F_{\alpha}$ for any ${\textsl{g}}\in{\rm SO}(3)$, thus (4.11) corresponds with the ${\rm SO}(5)$-invariant curvature $2$-form of the induced by the ${\rm SO}(5)$-invariant connection $\alpha\in\Omega^{1}(P,\mathfrak{so}(3))$ (cf. [cap2009, Proposition 1.4.6]).
4.2 Homogeneous $\mathrm{G}_{2}$ and self dual contact instantons (SDCI)
For the ${\rm SO}(5)$-invariant contact form $\eta=X^{1}$ on $V^{5,2}$ and the ${\rm SO}(5)$-invariant metric given by (2.7), we consider the equation
$\ast_{g}(F_{\beta}\wedge\eta\wedge\omega)=\pm F_{\beta},\quad\text{where}\quad\omega=\frac{1}{2}d\eta=(X^{25}+X^{36}+X^{47}).$ (4.12)
From Lemma 4.6, we obtain the following
Corollary 4.7.
Consider $V^{5,2}$ with the Sasakian structure given in Theorem 2.2. Then, the connection (4.8) is an invariant SDCI on the homogeneous ${\rm SO}(3)$–principal bundle $P_{1}\to V^{5,2}$.
Proof.
For any Sasakian manifold, the bundle of $2$–forms decomposes into [portilla2023, Section 2.1]
$\Omega^{2}=\Omega^{2}_{1}\oplus\Omega^{2}_{6}\oplus\Omega^{2}_{8}\oplus\Omega^{2}_{V},$ (4.13)
where (3.3) are elements of $\Omega^{2}_{8}=span\{w_{i}\}_{i=1}^{8}$ solving $\ast(\eta\wedge\omega\wedge w_{i})=w_{i}$. Comparing (4.11) with the values of $w_{i}$ in (3.3), we note that $F_{\alpha}$ in (4.11) is given by
$F_{\alpha}:=\frac{1}{y_{2}}(w_{1}\otimes e_{8}+w_{3}\otimes e_{9}+w_{5}\otimes e_{10}).$
The curvature tensor $F_{\alpha}$ satisfies $\ast(F_{\alpha}\wedge\eta\wedge\omega)=F_{\alpha}$. Hence, for any $y_{2}\in\mathbb{R}^{+}$ satisfying (2.22), the connection (4.8) is an ${\rm SO}(5)$-invariant SDCI. ∎
Corollary 4.8.
Consider $V^{5,2}$ endowed with the Sasakian structure given in Corollary 2.8. There exist an unique invariant ASDCI on the trivial homogeneous ${\rm SO}(3)$–principal bundle $P_{0}\to V^{5,2}$ (cf. Lemma 4.4).
Proof.
From Lemma 4.4, we only should note that the curvature (4.4) satisfy $\ast_{g}(F_{\beta}\wedge\eta\wedge\omega)=F_{\beta},$ if and only if $y_{1}=\frac{9}{2}$. ∎
A result analogous to Corollary 4.8 follows for $P\to V^{5,2}$ in Remark 4.5, since the curvature tensor in (4.6) satisfy $\ast_{g}(F_{\beta}\wedge\eta\wedge\omega)=F_{\beta},$ if and only if $y_{1}=\frac{9}{2}$.
Proposition 4.9.
Consider the connection (4.8) on $V^{5,2}$, with respect to the $\mathrm{G}_{2}$–structure given in (2.11). Then we gave
  • The ${\rm SO}(3)$–connection in (4.8) on the homogeneous principal bundle $P_{1}\to V^{5,2}$ [cf. Lemma 4.6] satisfies $F_{\alpha}\wedge\psi=0$ for any $y_{1}>0$, i.e., there exist an unique ${\rm SO}(5)$–invariant connection which is simultaneously $\mathrm{G}_{2}$ and SDCI, given by (4.11).
  • The ${\rm SO}(3)$–connection (4.1) on the trivial bundle $P_{0}\to V^{5,2}$ (cf. Lemma 4.4), whose curvature is given in (4.4) satisfies $F_{\alpha}\wedge\psi=0$, for $\psi$ given in (2.15), if and only if (4.1) is flat.
4.3 Stability of SDCI and $\mathrm{G}_{2}$-instantons
Now, we analyse the Yang-Mills condition for instanton (4.8), for a detailed study on stability of Yang-Mills connections we refer to [bourguignon1981stability]. On the space of connections $\mathcal{A}$, the global inner product $(\alpha,\beta)=\int_{M}\langle\alpha\wedge\ast\beta\rangle$ allow us define the Yang-Mills functional as the $L^{2}$-norm of the curvature:
$\mathcal{YM}\colon\nabla\in\mathcal{A}\to\|{\rm F}_{\nabla}\|^{2}_{L^{2}}=\int_{M}\operatorname{Tr}({\rm F}_{\nabla}\wedge\ast{\rm F}_{\nabla})=\int_{M}\|{\rm F}_{\nabla}\|^{2}\in\mathbb{R}.$ (4.14)
A critical point of (4.14) is called a Yang-Mills (YM) connection, i.e. $\nabla$ satisfies the so-called Yang-Mills equation $d^{\ast}_{\nabla}F_{\nabla}=0$. Taking the exterior derivative $d_{\nabla}$ in (1.2) and using the Bianchi identity, one note that a solution of the instanton equation (1.2) satisfy the YM-equation with torsion:
$d_{\nabla}(\ast F_{\nabla})=d\sigma\wedge F_{\nabla}.$
Note that neither $\sigma=\eta\wedge d\eta$ nor $\sigma=\varphi$ are closed, consequently, the torsion term in the above equation is no automatically zero, however, we have that SDCI and ASDCI are Yang-Mills connections [luis2024Weitzenbock, Section 5]. With respect to the $\mathrm{G}_{2}$-structure (2.23), we have the following:
Theorem 4.1.
Let $\alpha$ be the unique $\mathrm{G}_{2}$-instanton (4.8) on the homogeneous ${\rm SO}(3)$-principal bundle $P_{1}$. Then, $\alpha$ is a critical point for the Yang-Mills energy for any coclosed $\mathrm{G}_{2}$-structure (2.23).
Proof.
Using (2.17) and the fact that $\alpha$ is a $\mathrm{G}_{2}$-instanton, we have that the curvature (4.11) satisfies $d_{\alpha}^{*}F_{\alpha}=0$, if and only if
$2a^{3}-2a^{2}x-3ab^{2}+2aby+bxy=0,\quad\text{and}\quad 2b^{3}-2b^{2}y-3a^{2}b+2abx+axy=0,$
we see that both equations hold for (2.23), i.e. $x=a$ and $y=b$. Hence, we obtain the result. ∎
For the Sasakian case, we can prove that SDCI are local minimizers of the YM-functional (4.14). In order to prove this, note that the contact structure induces an orthogonal decomposition (cf. [portilla2023, Lemma 2.3]) of the curvature $2$-form ${\rm F}_{\nabla}$ into
${\rm F}_{H}=i_{\xi}(\eta\wedge{\rm F}_{\nabla})\quad\text{and}\quad{\rm F}_{V}=\eta\wedge i_{\xi}{\rm F}_{\nabla},$
where $F_{H}$ is the horizontal part and $F_{V}$ the vertical part of $F_{\nabla}$. Consequently, we have $\|{\rm F}_{\nabla}\|^{2}=\|{\rm F}_{H}\|^{2}+\|{\rm F}_{V}\|^{2}$ and denote
${\rm F}_{H}^{+}=\frac{1}{2}({\rm F}_{H}+\ast(\sigma\wedge{\rm F}_{H}))\quad\text{and}\quad{\rm F}_{H}^{-}=\frac{1}{2}({\rm F}_{H}-\ast(\sigma\wedge{\rm F}_{H})),$ (4.15)
According to the splitting (3.2), we have ${\rm F}_{H}={\rm F}_{8}+{\rm F}_{6}+{\rm F}_{1}\in\Omega_{8}^{2}({\mathfrak{g}}_{E})\oplus\Omega_{6}^{2}({\mathfrak{g}}_{E})\oplus\Omega_{1}^{2}({\mathfrak{g}}_{E})$, then we observe that ${\rm F}_{H}^{+}={\rm F}_{8}-\frac{1}{2}{\rm F}_{1}$ and ${\rm F}_{H}^{-}={\rm F}_{6}+\frac{3}{2}{\rm F}_{1}$, hence
$\|{\rm F}_{H}\|^{2}$ $=\langle{\rm F}_{H}^{+}+{\rm F}_{H}^{-},{\rm F}_{H}^{+}+{\rm F}_{H}^{-}\rangle=\|{\rm F}_{H}^{+}\|^{2}+\|{\rm F}_{H}^{-}\|^{2}+2\langle{\rm F}_{H}^{+},{\rm F}_{H}^{-}\rangle$
$=\|{\rm F}_{H}^{+}\|^{2}+\|{\rm F}_{H}^{-}\|^{2}-\frac{3}{2}\|{\rm F}_{1}\|^{2},$
consequently, the YM-functional (4.14) splits as
$\mathcal{YM}(\nabla)=\|{\rm F}_{H}^{+}\|^{2}+\|{\rm F}_{H}^{-}\|^{2}+\|{\rm F}_{V}\|^{2}-\frac{3}{2}\|{\rm F}_{1}\|^{2}.$ (4.16)
In the Sasakian context, the charge of the connection $\nabla$ is defined by
$\kappa(\nabla)=\displaystyle\int_{M}\mathop{\mathrm{Tr}}\nolimits({\rm F}_{\nabla}^{2})\wedge\eta\wedge\omega.$ (4.17)
Notice that the charge (4.17) is determined by the horizontal part of $F_{\nabla}$ [portilla2023, Proposition 3.13]:
$\kappa(\nabla)=\|{\rm F}_{H}^{+}\|^{2}-\|{\rm F}_{H}^{-}\|^{2}.$ (4.18)
Using (4.16) and (4.18), we have that the YM-mills action is given by
$\mathcal{YM}(\nabla)$ $=\kappa(\nabla)+2\|{\rm F}_{H}^{-}\|^{2}-\frac{3}{2}\|{\rm F}_{1}\|^{2}+\|{\rm F}_{V}\|^{2}$
$=-\kappa(\nabla)+2\|{\rm F}_{H}^{+}\|^{2}-\frac{3}{2}\|{\rm F}_{1}\|^{2}+\|{\rm F}_{V}\|^{2},$
hence, we have
$\mathcal{YM}(\nabla)=\begin{cases}\kappa(\nabla),&\nabla\in\Omega^{2}_{8}({\mathfrak{g}}_{E})\\ -\kappa(\nabla),&\nabla\in\Omega^{2}_{6}({\mathfrak{g}}_{E})\\ 2\kappa(\nabla),&\nabla\in\Omega^{2}_{1}({\mathfrak{g}}_{E}).\end{cases}$ (4.19)
We recall that if $M$ is a contact manifold, a $k$-form $\alpha\in\Omega^{k}(M)$ is called transversal form if $i_{\xi}\alpha=0$, we denote the space of transversal forms by $\Omega^{k}_{H}(M)$. Consequently, the space of $k$-forms splits as
$\Omega^{k}(M)=\Omega^{k}_{H}(M)\oplus\eta\wedge\Omega^{k-1}_{H}(M).$ (4.20)
More generally, consider a Riemannian manifold $M$ endowed with a Riemannian foliation $\mathcal{F}\subset M$ and consider a principal $G$-bundle $\pi\colon P\to M$. $P$ is said to be a foliated principal bundle if there exists a $G$-invariant foliation $\widetilde{\mathcal{F}}$ on $P$ such that the tangent distribution $T\widetilde{\mathcal{F}}\subset TP$ projects isomorphically onto $T\mathcal{F}$ under $\pi_{\ast}\colon TP\to TM$. Accordingly, a vector bundle is called foliated if its associated principal bundle is foliated. A local section $s\colon U\subset M\to P$ is said basic section if $s_{\ast}\colon TU\to TP$ maps $T\mathcal{F}$ to $T\widetilde{\mathcal{F}}$. It is known that every foliated principal bundle admits local basic sections [molino1988riemannian, Proposition 2.7]. We denote the space of basic sections by $\Gamma(E)_{B}\subset\Gamma(E)$. In the context of contact manifolds, $\mathcal{F}$ is naturally given by the Reeb foliation, if $\nabla\in\Gamma(E)$ is a connection with $i_{\xi}F_{\nabla}=0$, then the horizontal lift of $\mathcal{F}$ with respect to $\nabla$ is an integrable distribution and gives $E$ a foliated structure for which $\nabla$ is basic connection [molino1988riemannian], consequently, we have a natural split of the bundle of $E$-valued $k$-forms analogous to (4.20).
On the underlying bundle $E$, considering only transverse deformations, i.e., $\nabla^{\prime}=\nabla+\alpha$, where $\alpha\in\Omega^{1}_{H}({\mathfrak{g}}_{E})$, in [portilla2023, Lemma 3.14] is shown that the charge vary by
$\kappa(\nabla)=\kappa(\nabla^{\prime})+\int_{M}\mathop{\mathrm{Tr}}\nolimits({\rm F}_{\nabla}\wedge\alpha)\wedge\omega^{2}.$ (4.21)
The following result emulates [portilla2023, Lemma 3.14], the proof proceeds with the same reasoning, but we emphasise that, first, this result is not exclusive to the Stiefel manifold, but rather applies to any Sasakian manifold. And secondly, in [portilla2023] the result is announced for transverse deformations of the Chern connection [portilla2023, Corollary 3.15], but actually this result holds in a more general context as follow:
Lemma 4.10.
Let $E$ be a Sasakian ${\rm G}$-bundle over a compact closed 7-manifold. For any variation $\nabla^{\prime}=\nabla+\alpha\in\mathcal{A}(E)$, $\alpha\in\Omega^{1}({\mathfrak{g}}_{E})$, on the underlying complex vector bundle E, we have
$\kappa(\nabla)=\kappa(\nabla^{\prime})+\int_{M}\mathop{\mathrm{Tr}}\nolimits({\rm F}_{\nabla}\wedge\ast(\omega\otimes s)),$ (4.22)
where $s\in\Gamma(E)_{B}$ is some transversal section of $E$.
Proof.
The proof of [portilla2023, Lemma 3.14] provide that for an arbitrary deformation $\nabla^{\prime}=\nabla+\alpha$, where $\alpha\in\Omega^{1}({\mathfrak{g}}_{E})=\Omega^{1}_{H}({\mathfrak{g}}_{E})\oplus\eta\wedge\Omega^{0}_{H}({\mathfrak{g}}_{E})$, the charge varies as
$\kappa(\nabla)$ $=\kappa(\nabla^{\prime})+\int_{M}\mathop{\mathrm{Tr}}\nolimits({\rm F}_{\nabla}\wedge\alpha)\wedge\omega^{2}$ (4.23)
$\quad+\frac{1}{2}\int_{M}\mathop{\mathrm{Tr}}\nolimits(d_{\nabla}\alpha\wedge\alpha)\wedge\omega^{2}+\frac{1}{3}\int_{M}\mathop{\mathrm{Tr}}\nolimits(\alpha\wedge\alpha\wedge\alpha)\wedge\omega^{2}.$
In particular, assuming a transversal deformation, (4.21) holds because the defect term in (4.23) vanishes, as the subintegral quantities in the last two terms belong to $\Omega^{7}_{H}({\mathfrak{g}}_{E})=0$. Hence, we can assume that $\alpha$ is a vertical $1$-form, i.e., $\alpha=\eta\otimes s\in\eta\wedge\Omega^{0}_{H}({\mathfrak{g}}_{E})$ for some transversal section $s$, therefore
$\kappa(\nabla)$ $=\kappa(\nabla^{\prime})+\int_{M}\mathop{\mathrm{Tr}}\nolimits({\rm F}_{\nabla}\wedge(\eta\otimes s))\wedge\omega^{2}+\frac{1}{2}\int_{M}\mathop{\mathrm{Tr}}\nolimits(d_{\nabla}(\eta\otimes s)\wedge(\eta\otimes s))\wedge\omega^{2}$
$=\kappa(\nabla^{\prime})+\int_{M}\mathop{\mathrm{Tr}}\nolimits({\rm F}_{\nabla}\wedge\eta\wedge\omega^{2}\otimes s)+\frac{1}{2}\int_{M}\mathop{\mathrm{Tr}}\nolimits(\omega\otimes s-\eta\otimes\nabla s)\wedge(\eta\otimes s))\wedge\omega^{2}$
$=\kappa(\nabla^{\prime})+\int_{M}\mathop{\mathrm{Tr}}\nolimits({\rm F}_{\nabla}\wedge\ast(\omega\otimes s))+\frac{1}{2}\int_{M}\mathop{\mathrm{Tr}}\nolimits(\eta\wedge\omega^{3}\otimes[s,s]),$
hence, (4.22) follows. ∎
Corollary 4.11.
Let $E$ be a Sasakian ${\rm G}$-bundle over a compact closed 7-manifold. Any SDCI is a local minimizer of the YM-functional (4.14).
Proof.
From (4.19), note that the YM-functional assumes the value of the charge $\kappa(\nabla)$ (4.17). From (4.22)
$\kappa(\nabla)$ $=\kappa(\nabla^{\prime})+({\rm F}_{\nabla},\omega\otimes s)_{L^{2}}$
$=\kappa(\nabla^{\prime}),$
the last line follows since $\Omega^{2}_{8}({\mathfrak{g}}_{E})\perp\omega\otimes\Omega^{0}({\mathfrak{g}}_{E})$. Hence, if $\nabla$ is a SDCI the charge is locally constant. The same is true if $\nabla$ is a ASDCI, since $\Omega^{2}_{6}({\mathfrak{g}}_{E})\perp\omega\otimes\Omega^{0}({\mathfrak{g}}_{E})$. ∎