### v2

# Homogeneous $\mathrm{G}_{2}$ and Sasakian instantons on the Stiefel $7$-manifold

Abstract
We study homogeneous instantons on the seven–dimensional Stiefel manifold $V^{5,2}$ in the context of $\mathrm{G}_{2}$ and Sasakian geometry. According to the reductive decomposition of $V^{5,2}$, we provide an explicit description of all invariant $\mathrm{G}_{2}$ and Sasakian structures. In particular, we characterise the invariant $\mathrm{G}_{2}$–structures inducing a Sasakian metric, among which the well–known nearly parallel $\mathrm{G}_{2}$–structure (Sasaki–Einstein) is included. As a consequence, we classify the invariant connections on homogeneous principal bundles over $V^{5,2}$ with gauge group ${\rm U}(1)$ and ${\rm SO}(3)$, satisfying either the $\mathrm{G}_{2}$ or the Sasakian instanton condition. In addition, we study infinitesimal deformations of $\mathrm{G}_{2}$–instantons on coclosed $\mathrm{G}_{2}$–manifolds using a spinorial approach. By means of a Weitzenböck–type formula with torsion, we obtain curvature obstructions to the existence of non–trivial infinitesimal deformations and prove rigidity results for certain homogeneous $\mathrm{G}_{2}$–instantons.
1 Introduction
The Stiefel manifold $V^{n,k}$ of orthonormal $k$-frames in the $n$-dimensional Euclidean space (i.e. $1\leq k\leq n$) admits a transitive ${\rm SO}(n)$-action with isotropy subgroup isomorphic to ${\rm SO}(n-k)$, thus, it can be written homogeneously as $V^{n,k}={\rm SO}(n)/{\rm SO}(n-k)$ and consequently, its family of ${\rm SO}(n)$-invariant Riemannian metrics is described in terms of the isotropy representation of $V^{n,k}$. For instance, for $k=1$ or $k=n$, the isotropy representation is irreducible (indeed $V^{n,1}=S^{n}$ or $V^{n,n}={\rm SO}(n)$), in those cases, there is a unique ${\rm SO}(n)$-invariant metric, which is Einstein [wolf1968]. For $k=2$, the isotropy representation is reducible and the space of ${\rm SO}(n)$-invariant Riemannian metrics is described by a $4$-dimensional family, within which there exists a unique Einstein metric, which is inherited from the Grassmannian $G_{2}(\mathbb{R}^{n})$ (e.g. [Arvanitoyeorgos1996, kerr1998new]). In addition, $V^{n,2}$ fits as an instance of the Boothby-Wang construction, as the total space of a $S^{1}$-principal bundle [boothby1958]*Theorem 3, moreover, it has a regular contact structure whose associated Reeb vector field generates a one-parameter subgroup of right translations of $V^{n,2}$.
A contact structure [^1] The reader unfamiliar with Sasakian geometry can find a complete study on this subject in [Boyer2008] on a $(2n+1)$-dimensional manifold $M$ is a 1-form satisfying $\eta\wedge(d\eta)^{n}\neq 0$, in this case there exists a unique vector $\xi\in\mathscr{X}(M)$, called Reeb vector field, such that $\eta(\xi)=1$. The contact structure is called an almost contact structure if there exists an endomorphism $\Phi\in{\mathrm{End}}(TM)$ which satisfy $\Phi^{2}=-{\rm I}_{TM}+\eta\otimes\xi.$ A Riemannian metric $g$ is called compatible with the contact structure, if $g$ satisfies $g(\Phi X,\Phi Y)=g(X,Y)-\eta(X)\eta(Y),$ in this case, the contact structure is called an almost contact metric structure. In addition, if $\mathcal{S}=(\eta,\xi,g,\Phi)$ satisfies $2g(X,\Phi(Y))=d\eta(X,Y)$, the quadruple $\mathcal{S}$ is called a contact metric structure. Furthermore, if the Nijenhuis tensor $N_{\Phi}$ induced by $\Phi$ satisfies $N_{\Phi}=-d\eta\otimes\xi$, then $\mathcal{S}$ is called a Sasakian structure. In this paper, we consider the 77–dimensional Stiefel manifold $V^{5,2}$ endowed with a one-parameter family of ${\rm SO}(5)$-invariant Sasakian structures $\mathcal{S}=(\eta,\xi,g,\Phi)$. For instance, homogeneous contact/Sasakian structures have been studied in [andrada2009, correa2019] and [perrone1998homogeneous].
On the other hand, a differential $3$-form $\varphi$ on an oriented 77–dimensional manifold $M$ is called a $\mathrm{G}_{2}$–structure if at each $p\in M$, there is a basis $e_{1},\dots,e_{7}$ of $T_{p}M$ such that
$\varphi_{p}=e^{125}+e^{136}+e^{147}+e^{234}-e^{267}+e^{357}-e^{456},$ (1.1)
where $e^{1},\dots,e^{7}$ denotes the dual basis of $T_{p}^{\ast}M$ and $e^{ijk}=e^{i}\wedge e^{j}\wedge e^{k}$. In particular, $\varphi$ determines a Riemannian metric $g_{\varphi}$, we study the family of ${\rm SO}(5)$-invariant $\mathrm{G}_{2}$-structures on $V^{5,2}$, such that the induced metric $g_{\varphi}$ is Sasakian. With respect to these structures, we study the theoretical gauge instanton equations associated with the underlying Sasakian and $\mathrm{G}_{2}$-geometry of $V^{5,2}$. Namely, on a GG–principal bundle $P\to M^{7}$ with $\sigma\in\Omega^{3}(M)$, the $1$-form connection $A\in\mathcal{A}(P)$ is called an $\lambda$-instanton, if it satisfies (e.g. [donaldson1998gauge, Tian2000])
$\ast(\sigma\wedge F_{A})=\lambda F_{A},\quad\text{where}\quad\lambda\in\mathbb{R}$ (1.2)
where $F_{A}=dA+A\wedge A$ denotes the corresponding curvature $2$-form. According to the specific pair $(\sigma,\lambda)$ in (1.2), we define:
  1. 1.
    A Self-dual contact instantons (SDCI) [portilla2023] is a solution of (1.2) with $\lambda=1$ and $\sigma:=\eta\wedge\omega$ naturally defined by a contact structure of $(M,\eta)$, where $\omega:=\frac{1}{2}d\eta$ is called the fundamental $2$-form.
  2. 2.
    A $\mathrm{G}_{2}$–instantons [corrigan1983first, donaldson1998gauge] is a solution of (1.2) with $\lambda=-1$ and $\sigma=\varphi$ a coclosed (i.e. $d\ast\varphi=0$) $\mathrm{G}_{2}$–structure on $M$. Equivalently, a connection $A$ is a $\mathfrak{g}$}–instanton if and only if
    $F_{A}\wedge\ast\varphi=0$ (1.3)
Furthermore, for $E\to(M^{7},\mathcal{S})$ a Sasakian holomorphic bundle [biswas2010vector], the usual notion of Hermitian Yang-Mills (HYM) connection on a Kähler manifolds is extended to odd dimension in the natural fashion to the so-called transverse Hermitian Yang-Mills connections (tHYM) (see (3.1)). The concept of Chern connections also extends naturally to a Sasakian version, as a connection mutually compatible with the transverse holomorphic structure and the Hermitian metric. A noteworthy result is that if $E\to M$ is a holomorphic Sasakian bundle on a contact Calabi-Yau manifold $M$, along compatible connections the notion of SDCI, $\mathrm{G}_{2}$-instantons and tHYM coincide. cf. [portilla2023, Theorem 1.1] and [calvo2020gauge, Lemma  21].
In [donaldson1998gauge, Donaldson2011], Donaldson, Segal and Thomas suggested that it may be possible to use $\mathrm{G}_{2}$-instantons to construct an enumerative invariant of $\mathrm{G}_{2}$-manifolds. For instance, when $(M^{7},\varphi)$ has a torsion-free $\mathrm{G}_{2}$-structure $\varphi$ (i.e. $d\varphi=0$ and $d\ast\varphi=0$ [fernandez1982riemannian]), solutions to the $\mathrm{G}_{2}$-instanton equation have been studied in [sa2015g2, walpuski2015g_2, clarke2014instantons, lotay20182, singhal2022]. However, the construction of metrics with holonomy $\mathrm{G}_{2}$ has proven to be a challenging task, thus, the torsion-free condition has been relaxed to the class of coclosed $\mathrm{G}_{2}$-structures [ball2019gauge, clarke2023], which has a notable application in theoretical physics to the Killing spinor equations in supergravity (cf. [Ivanov2005, Ivanov10, delaOssaG21, Lotay2022, Fernandez2011]). Analogously, Wang has proposed studying the compactification of the moduli space of contact instantons as a tool to construct a Donaldson-type invariant for contact $7$-manifolds [wang2012higher].
Description of the main results: Let $\mathfrak{so}(5)=\mathfrak{so}(3)\oplus{\mathfrak{m}}$ be a reductive decomposition of $V^{5,2}$, in [reidegeld2010spaces] Reidegeld proved that the space of ${\rm SO}(5)$-invariant $\mathrm{G}_{2}$-structures has dimension $5$ and the space of induced ${\rm SO}(5)$-invariant Riemannian metrics has dimension $4$. Thus, we prove in Lemma 2.3 that the $\mathrm{Ad}({\rm SO}(3))$-invariant $3$-form (2.11) determined by $(a,b,x,y,z)\in\mathbb{R}^{5}$ defines a $\mathrm{G}_{2}$-structure (up to $\mathrm{G}_{2}$-transformation), if
$a^{2}+by>0,\quad b^{2}+ax>0\quad z\neq 0\quad\text{and}\quad xy=ab.$
In particular, it defines a family of (non-purely) coclosed $\mathrm{G}_{2}$-structures, within the nearly parallel $\mathrm{G}_{2}$-structures arises as an $S^{1}$–family for $x=a$, $y=b$ and $a^{2}+b^{2}=z^{2}=\frac{27}{512}$ (cf. Theorem 2.1).
In addition, for any $y_{1},y_{2},y_{3}\in\mathbb{R}^{+}$, Theorem 2.2 characterises the ${\rm SO}(5)$-invariant Sasakian metrics as
$y_{1}=4y_{2}^{2}\quad\text{and}\quad y_{2}=y_{3},$
as a consequence, for a given $(z,a,b,x,y)\in\mathbb{R}^{5}$, we describe the family of ${\rm SO}(5)$-invariant $\mathrm{G}_{2}$-structures inducing a Sasakian metric (cf. Corollary 2.8).
In Section 3, we revisit the interaction between SDCI, $\mathrm{G}_{2}$-instantons and tHYM connections addressed in [portilla2023, Proposition 3.11] and [calvo2020gauge, Lemma 21] against the backdrop of contact Calabi-Yau manifolds. For a bundle $E\to V^{5,2}$, in Lemma 3.1 we prove that if the connection $A$ is SDCI then it is a $\mathrm{G}_{2}$-instanton. Conversely, for a holomorphic Sasakian bundle $E\to V^{5,2}$, if the Chern connection is a $\mathrm{G}_{2}$-instanton, then it is a SCDI. Finally, we prove that HYM connections on a holomorphic bundle $E\to G_{2}(\mathbb{R}^{5})$ on the Grassmannian, pulling back to $\mathrm{G}_{2}$-instantons on $\pi^{\ast}E\to V^{5,2}$ (cf. Corollary 3.4). We point out that the key argument can be carried out under the sole assumption of an $\mathrm{SU}(3)$-structure, thereby improving the results mentioned above.
In Section 4, we provide examples of homogeneous SDCI and $\mathrm{G}_{2}$-instantons over $V^{5,2}$, with gauge group ${\rm SO}(3)$ and ${\rm U}(1)$. In particular, we prove that the ${\rm SO}(3)$-invariant $\mathrm{G}_{2}$-instanton (4.9) is a Yang-Mills connection on $(V^{5,2},\varphi)$, for any ${\rm SO}(5)$-invariant coclosed $\mathrm{G}_{2}$-structure (cf. Theorem 4.1).
Finally, in Section 5 we extend the spinorial framework proposed by R. Singhal [singhal2022] to coclosed $\mathrm{G}_{2}$–structures, and as a concrete application in Corollary 5.7, we prove rigidity for a family of homogeneous $\mathrm{G}_{2}$–instantons on the Stiefel manifold $V^{5,2}$ for a non-empty range of the metric parameter $y_{2}$.
Outline: In Section 2, we provide some preliminaries about $V^{5,2}$ as a homogeneous space, with the aim of characterising all the ${\rm SO}(5)$-invariant coclosed $\mathrm{G}_{2}$-structures and the associated Sasakian metrics. In Section 3, we explore the interaction between SDCI and $\mathrm{G}_{2}$–instantons on $V^{5,2}$, in particular, Lemmata 3.1, 3.3 and Corollary 3.4 indicate conditions such that contact instantons and $\mathrm{G}_{2}$-instantons coincide. Finally, in Section 4 we consider homogeneous principal bundles on $V^{5,2}$ and we construct examples of invariant instantons on ${\rm SO}(3)$, U(1)U(1)–homogeneous bundles. In particular, in Section 4.1 we characterise all invariant ${\rm SO}(3)$–connection and we show that these satisfy the $\mathrm{G}_{2}$ and SDCI condition for appropriate parameters. Finally, in Section 5 we study the deformation theory for a particular example of a $\mathrm{G}_{2}$-instanton, specifically in Corollary 5.7 we prove that it is rigid for a non-empty interval of the metric parameter $y_{2}$.
Acknowledgments and funding: AM was funded by the São Paulo Research Foundation [2021/08026-5] and [2023/13780-6] (Fapesp). LP was funded by Centre Henri Lebesgue (CHL) under the grant ANR-11-LABX-0020-01. He also thanks the LMBA at Université de Bretagne Occidentale (UBO) for its hospitality
2 Homogeneous $\mathrm{G}_{2}$ and contact metric structures on $V^{5,2}$
In this section, we provide an explicit description of a $\mathrm{Ad}({\rm SO}(3))$-invariant $\mathrm{G}_{2}$-structure and the induced inner product on $V^{5,2}$. We also establish the notation that we will use throughout this work.
2.1 $\mathrm{G}_{2}$-structures on $V^{5,2}$
In order to describe the space of ${\rm SO}(5)$-invariant $\mathrm{G}_{2}$-structures on $V^{5,2}={\rm SO}(5)/{\rm SO}(3)$ [^2] This subject is discussed in depth in [friedrich1997nearly]., we fix a reductive decomposition of $V^{5,2}$ as a homogeneous space. Consider the Lie algebra of ${\rm SO}(5)$
$\mathfrak{so}(5):=\{A\in\mathfrak{gl}(\mathbb{R}^{5})\;|\;A+A^{T}=0\},$ (2.1)
Let $E_{ij}$ denote the elementary $5\times 5$ matrix with a $1$ in the $(i,j)$-entry and zeros elsewhere. We fix the following basis of the Lie algebra $\mathfrak{so}(5)$:
$\begin{split}e_{1}=&E_{12}-E_{21},\quad e_{4}=E_{15}-E_{51},\quad e_{7}=E_{25}-E_{52},\quad e_{10}=E_{54}-E_{45},\\[3.00003pt] e_{2}=&E_{13}-E_{31},\quad e_{5}=E_{23}-E_{32},\quad e_{8}=E_{34}-E_{43},\quad\\[3.00003pt] e_{3}=&E_{14}-E_{41},\quad e_{6}=E_{24}-E_{42},\quad e_{9}=E_{35}-E_{53}.\end{split}$ (2.2)
We consider the Lie algebra $\mathfrak{so}(3)$ embedded in $\mathfrak{so}(5)$ which is spanned by $\{e_{8},e_{9},e_{10}\}$. The Lie bracket of $\mathfrak{so}(5)$ is given by the commutator described in Table 1.
$\cdot$$\cdot$] $e_{1}$ $e_{2}$ $e_{3}$ $e_{4}$ $e_{5}$ $e_{6}$ $e_{7}$ $e_{8}$ $e_{9}$ $e_{10}$
$e_{1}$ 0 $e_{5}$ $e_{6}$ $e_{7}$ $e_{2}$ $e_{3}$ $e_{4}$ 0 0 0
$e_{2}$ $e_{5}$ 0 $e_{8}$ $e_{9}$ $e_{1}$ 0 0 $e_{3}$ $e_{4}$ 0
$e_{3}$ $e_{6}$ $e_{8}$ 0 $e_{10}$ 0 $e_{1}$ 0 $e_{2}$ 0 $e_{4}$
$e_{4}$ $e_{7}$ $e_{9}$ $e_{10}$ 0 0 0 $e_{1}$ 0 $e_{2}$ $e_{3}$
$e_{5}$ $e_{2}$ $e_{1}$ 0 0 0 $e_{8}$ $e_{9}$ $e_{6}$ $e_{7}$ 0
$e_{6}$ $e_{3}$ 0 $e_{1}$ 0 $e_{8}$ 0 $e_{10}$ $e_{5}$ 0 $e_{7}$
$e_{7}$ $e_{4}$ 0 0 $e_{1}$ $e_{9}$ $e_{10}$ 0 0 $e_{5}$ $e_{6}$
$e_{8}$ 0 $e_{3}$ $e_{2}$ 0 $e_{6}$ $e_{5}$ 0 0 $e_{10}$ $e_{9}$
$e_{9}$ 0 $e_{4}$ 0 $e_{2}$ $e_{7}$ 0 $e_{5}$ $e_{10}$ 0 $e_{8}$
$e_{10}$ 0 0 $e_{4}$ $e_{3}$ 0 $e_{7}$ $e_{6}$ $e_{9}$ $e_{8}$ 0
Table 1: Bracket of $\mathfrak{so}(5)$ with respect to the basis (2.2)
On $\mathfrak{so}(5)$, we fix the inner product $\langle a,b\rangle=Q(a,b)=\frac{1}{2}\mathop{\mathrm{Tr}}\nolimits(ab^{T}),$ with respect to which, we have the orthogonal splitting
$\mathfrak{so}(5)={\mathfrak{m}}\oplus\mathfrak{so}(3)\quad\text{where}\quad{\mathfrak{m}}:=\mathfrak{so}(3)^{\perp}=\operatorname{{\rm Span}}\{e_{1},...,e_{7}\}.$ (2.3)
Note that in terms of the basis (2.2), the restriction of $\mathrm{ad}(e_{k})_{{\mathfrak{m}}}$ to $\mathfrak{m}$} are given by
$\mathrm{ad}(e_{8})_{\mathfrak{m}}$ $=\mathrm{diag}(0,A_{1},A_{1}),$
$\mathrm{ad}(e_{9})_{\mathfrak{m}}$ $=\mathrm{diag}(0,A_{2},A_{2}),$
$\mathrm{ad}(e_{10})_{\mathfrak{m}}$ $=\mathrm{diag}(0,A_{3},A_{3}),$
where $A_{1}=\begin{pmatrix}0&-1&0\\ 1&0&0\\ 0&0&0\end{pmatrix}$,   $A_{2}=\begin{pmatrix}0&0&-1\\ 0&0&0\\ 1&0&0\end{pmatrix}$ and $A_{3}=\begin{pmatrix}0&0&0\\ 0&0&-1\\ 0&1&0\end{pmatrix}$ are the canonical basis of $\mathfrak{so}(3)$. Denote by ${\mathfrak{m}}_{1}=\operatorname{{\rm Span}}\{e_{1}\},$ ${\mathfrak{m}}_{2}:=\operatorname{{\rm Span}}\{e_{2},e_{3},e_{4}\}$ and ${\mathfrak{m}}_{3}:=\operatorname{{\rm Span}}\{e_{5},e_{6},e_{7}\}.$ By using the Lie bracket in Table 1, it is straightforward to verify that
$\mathrm{ad}(\mathfrak{so}(3)){\mathfrak{m}}_{1}=0,\quad\mathrm{ad}(\mathfrak{so}(3)){\mathfrak{m}}_{2}\subseteq{\mathfrak{m}}_{2}\quad\text{and}\quad\mathrm{ad}(\mathfrak{so}(3)){\mathfrak{m}}_{3}\subseteq{\mathfrak{m}}_{3}.$
Consequently, for the reductive decomposition (2.3) we obtain the irreducible $\mathrm{ad}(\mathfrak{so}(3))$-submodules
${\mathfrak{m}}={\mathfrak{m}}_{1}\oplus{\mathfrak{m}}_{2}\oplus{\mathfrak{m}}_{3}.$ (2.4)
The subalgebra $\mathfrak{so}(2)\oplus\mathfrak{so}(3)=\operatorname{{\rm Span}}\{e_{1},e_{8},e_{9},e_{10}\}\subset\mathfrak{so}(5)$ provides a reductive decomposition $\mathfrak{so}(5)=\mathfrak{so}(2)\oplus\mathfrak{so}(3)\oplus{\mathfrak{p}}$ (where ${\mathfrak{p}}={\mathfrak{m}}_{2}\oplus{\mathfrak{m}}_{3}$) for the symmetric space
$G_{5,2}=\frac{{\rm SO}(5)}{{\rm SO}(2)\times{\rm SO}(3)}.$ (2.5)
Furthermore, ${\mathfrak{p}}\simeq\mathbb{C}^{3}$ with respect to the canonical almost complex structure $J=\begin{pmatrix}0&-I_{3}\\ I_{3}&0\end{pmatrix}$ and it turns out that ${\mathfrak{m}}_{3}=\operatorname{{\rm Span}}\{Je_{2},Je_{3},Je_{4}\}$. With respect to the basis (2.2), we define the alternating forms on $\mathfrak{p}$} given by
$\begin{array}[]{rl}\omega_{0}&=e^{25}+e^{36}+e^{47},\\[4.0pt] \mathop{\mathrm{Im}}\Omega_{0}&=e^{237}-e^{246}+e^{345}-e^{567},\\[4.0pt] \mathop{\mathrm{Re}}\Omega_{0}&=e^{234}-e^{267}+e^{357}-e^{456}.\end{array}$ (2.6)
Notice that $\omega_{0}\wedge\mathop{\mathrm{Re}}\Omega_{0}=0$ and $2\omega_{0}^{3}=3\mathop{\mathrm{Re}}\Omega_{0}\wedge\mathop{\mathrm{Im}}\Omega_{0}\neq 0$, i.e. $(\omega_{0},\mathop{\mathrm{Re}}\Omega_{0})$ is a ${\rm SU}(3)$-structure on $\mathfrak{p}$} (with induced orientation $e^{234567}$).
Lemma 2.1.
The forms in (2.6) are $\mathrm{ad}(\mathfrak{so}(2)\oplus\mathfrak{so}(3))$-invariant, therefore, they induce a ${\rm SO}(5)$-invariant ${\rm SU}(3)$-structure $(\omega,\mathop{\mathrm{Re}}\Omega)$ on $G_{2}(\mathbb{R}^{5})$. Furthermore, we have that $d\omega=0,$ $d(\mathop{\mathrm{Re}}\Omega)=3e^{1}\wedge\mathop{\mathrm{Im}}\Omega$ and $d(\mathop{\mathrm{Im}}\Omega)=-3e^{1}\wedge\mathop{\mathrm{Re}}\Omega.$
Proof.
The Lie algebras $\mathfrak{sl}(\mathbb{C}^{3})$ and ${\mathfrak{s}}{\mathfrak{p}}(\mathbb{R}^{6})$ are described by:
$\mathfrak{sl}(\mathbb{C}^{3})=\left\{\begin{pmatrix}B&-C\\ C&B\end{pmatrix}\in\mathfrak{gl}({\mathfrak{p}})\Big|B,C\in\mathfrak{sl}(\mathbb{R}^{3})\right\},\quad{\mathfrak{s}}{\mathfrak{p}}(\mathbb{R}^{6})=\left\{\begin{pmatrix}B&C\\ D&-B^{T}\end{pmatrix}\in\mathfrak{gl}({\mathfrak{p}})\Big|C,D\in sym(\mathbb{R}^{3})\right\}.$
The Lie algebras $\mathfrak{sl}(\mathbb{C}^{3})$ and ${\mathfrak{s}}{\mathfrak{p}}(\mathbb{R}^{6})$ can be interpreted as stabilisers
$\mathfrak{stab}_{\mathfrak{gl}({\mathfrak{p}})}(\omega_{0})=\mathfrak{sl}(\mathbb{C}^{3})$
and
$\mathfrak{stab}_{\mathfrak{gl}({\mathfrak{p}})}(\mathop{\mathrm{Re}}\Omega_{0})=\mathfrak{stab}_{\mathfrak{gl}({\mathfrak{p}})}(\mathop{\mathrm{Im}}\Omega_{0})=\mathfrak{sp}(\mathbb{R}^{6}).$
Notice that the endomorphisms $\mathrm{ad}(e_{8})_{\mathfrak{p}}$, $\mathrm{ad}(e_{9})_{\mathfrak{p}}$, $\mathrm{ad}(e_{10})_{\mathfrak{p}}$ and $\mathrm{ad}(e_{1})_{\mathfrak{p}}=J$ belong to the intersection $\mathfrak{sl}(\mathbb{C}^{3})\cap\mathfrak{sp}(\mathbb{R}^{6})$. Therefore, the forms in (2.6) are $\mathrm{ad}(\mathfrak{so}(2)\oplus\mathfrak{so}(3))$-invariant. Now, for any $p=x\cdot\Big({\rm SO}(2)\times{\rm SO}(3)\Big)\in G_{5,2}$ define $(\omega,\real\Omega)$ by
$\omega_{p}={\rm L}_{x}^{\ast}\omega_{0}\quad\text{and}\quad\real\Omega_{p}={\rm L}_{x}^{\ast}\real\Omega_{0},$
where ${\rm L}_{x}$ is the left multiplication by $x\in{\rm SO}(5)$, hence $2\omega^{3}=3\mathop{\mathrm{Re}}\Omega\wedge\mathop{\mathrm{Im}}\Omega\neq 0$ . Finally, from Table 1 follows that $d\omega_{0}=0,$ $d(\mathop{\mathrm{Re}}\Omega_{0})=3e^{1}\wedge\mathop{\mathrm{Im}}\Omega_{0}$ and $d(\mathop{\mathrm{Im}}\Omega_{0})=-3e^{1}\wedge\mathop{\mathrm{Re}}\Omega_{0}$. ∎
The following result is mentioned in [kerr1998new]*§4. For the sake of completeness, we include the detailed proof.
Lemma 2.2.
Any $\mathrm{Ad}({\rm SO}(3))$-invariant metric on $V^{5,2}$ is isometric to a diagonal one, i.e., up to isometries, it is enough to consider the class of invariant metrics determined by the $\mathrm{Ad}({\rm SO}(3))$-invariant inner product:
$g=y_{1}Q|_{{\mathfrak{m}}_{1}}+y_{2}Q|_{{\mathfrak{m}}_{2}}+y_{3}Q|_{{\mathfrak{m}}_{3}},\quad\text{where}\quad y_{i}>0\quad i=1,2,3.$ (2.7)
Proof.
In the decomposition (2.4) we have ${\mathfrak{m}}_{2}\simeq{\mathfrak{m}}_{3}$, hence, any ${\rm SO}(5)$-invariant metric on $V^{5,2}$ is determined by the $\mathrm{Ad}({\rm SO}(3))$-invariant inner product $\langle\cdot,\cdot\rangle=Q(g\cdot,\cdot)$, where $g$ is a positive symmetric matrix of the form
$g=\begin{pmatrix}y_{1}&0&0\\ 0&y_{2}I_{3}&xI_{3}\\ 0&xI_{3}&y_{3}I_{3}\end{pmatrix}\quad\text{for}\quad y_{1},y_{2},y_{3}>0,\quad x\in\mathbb{R}\quad\text{with}\quad y_{2}y_{3}>x^{2}.$
Denote by $N({\rm SO}(3))$ the normalizer of ${\rm SO}(3)$ in ${\rm SO}(5)$, since the quotient $\frac{N({\rm SO}(3))}{{\rm SO}(3)}$ is isomorphic to ${\rm SO}(2)$ the conjugation induces a 1-parameter family of diffeomorphisms on $\mathfrak{m}$} given by $\phi(t)=\mathrm{Ad}(\exp(te_{1}))$. Thus, for $t\in\mathbb{R}$ the one-parameter family $g_{t}=\phi(t)^{T}g\phi(t)$ defines an isometric $\mathrm{Ad}({\rm SO}(3))$-inner product.
A direct computation shows that the parameter $x(t)$ in the metric $g_{t}$, is given by $x(t)=\frac{(y_{3}-y_{2})}{2}\sin(2t)+x\cos(2t)$. Hence, $g_{t}$ is diagonal if:
$y_{2}\neq y_{3}\quad\text{and}\quad t=\frac{1}{2}\arctan\left(\frac{2x}{y_{2}-y_{3}}\right)$
or
$y_{2}=y_{3}\quad\text{and}\quad t=\frac{(2k+1)\pi}{4}\quad\text{with}\quad k\in{\bf N},$
With respect to (2.7), we define an orthonormal basis of $\mathfrak{m}$}:
$X_{1}:=\frac{1}{\sqrt{y_{1}}}e_{1},\quad X_{i}:=\frac{1}{\sqrt{y_{2}}}e_{i}\quad\text{and}\quad X_{i+3}:=\frac{1}{\sqrt{y_{3}}}e_{i+3}\quad\text{for}\quad i=1,2,3.$ (2.8)
Then, the Lie bracket in Table 1 becomes
$\begin{split}[X_{1},X_{i}]_{{\mathfrak{m}}}=-\sqrt{\frac{y_{3}}{y_{1}y_{2}}}X_{i+3},\quad[X_{1},X_{i+3}]_{{\mathfrak{m}}}=\sqrt{\frac{y_{2}}{y_{1}y_{3}}}X_{i},\quad[X_{i},X_{i+3}]_{{\mathfrak{m}}}=-\sqrt{\frac{y_{1}}{y_{2}y_{3}}}X_{1}.\end{split}$ (2.9)
In [reidegeld2010spaces], Reidegeld proved that the space of ${\rm SO}(5)$-invariant $\mathrm{G}_{2}$-structures on $V^{5,2}$ has dimension $5$. The next Lemma describes the invariant $\mathrm{G}_{2}$-structures on the Stiefel manifold up to $\mathrm{G}_{2}$-equivalence, i.e. for a $\mathrm{G}_{2}$-structure $\varphi$ written in the basis $\{e_{1},...,e_{7}\}$ and $\{\tilde{e}_{1},...,\tilde{e}_{7}\}$ of $\mathfrak{p}$}, there is a change of basis $f\in\mathrm{G}_{2}$ from $\{e_{1},...,e_{7}\}$ to $\{\tilde{e}_{1},...,\tilde{e}_{7}\}$.
Lemma 2.3.
The $5$-dimensional family of ${\rm SO}(5)$-invariant $\mathrm{G}_{2}$-structures on $V^{5,2}$ is described, up to $\mathrm{G}_{2}$-equivalence, by $(a,b,x,y,z)\in\mathbb{R}^{5}$ satisfying
$a^{2}+by>0,\quad b^{2}+ax>0\quad z\neq 0\quad\text{and}\quad xy=ab.$ (2.10)
Explicitly, the $\mathrm{G}_{2}$-structure $\varphi$ can be written as
$\varphi=$ $-X^{1}\wedge(X^{25}+X^{36}+X^{47})+p_{1}X^{234}+p_{2}(-X^{267}+X^{357}-X^{456})$ (2.11)
$+p_{3}(X^{237}-X^{246}+X^{345})-p_{4}X^{567},$
where
$p_{1}=\frac{x\sqrt{a^{2}+by}}{b^{2}+ax},\quad p_{2}=\frac{a}{\sqrt{a^{2}+by}},\quad p_{3}=\frac{b}{\sqrt{b^{2}+ax}}\quad\text{and}\quad p_{4}=\frac{y\sqrt{b^{2}+ax}}{a^{2}+by}$ (2.12)
and
$y_{1}=\frac{z^{2}}{y_{2}y_{3}},\quad y_{2}=\frac{(b^{2}+ax)^{2/3}}{(a^{2}+by)^{1/3}}\quad\text{and}\quad y_{3}=\frac{(a^{2}+by)^{2/3}}{(b^{2}+ax)^{1/3}}.$ (2.13)
Proof.
From (2.6) it is straightforward to check that any $\mathrm{Ad}({\rm SO}(3))$-invariant $3$-form on $\mathfrak{m}$} is given by
$\widetilde{\varphi}$= $-\widetilde{z}e^{1}\wedge(e^{25}+e^{36}+e^{47})+\widetilde{x}e^{234}+\widetilde{a}(-e^{267}+e^{357}-e^{456})$
$+\widetilde{b}(e^{237}-e^{246}+e^{345})-\widetilde{y}e^{567}.$
The $3$-form $\widetilde{\varphi}$ defines a $\mathrm{G}_{2}$-structure if the induced symmetric bilinear form
$B_{ij}=\Big((e_{i}\lrcorner\widetilde{\varphi})\wedge(e_{j}\lrcorner\widetilde{\varphi})\wedge\widetilde{\varphi}\Big)(e_{1},...,e_{7})$
is positive (or negative) definite. Explicitly, we obtain that $B_{11}=6\widetilde{z}^{3}$ and for $k=2,3,4$
[…]
and $B_{ij}=0$ otherwise.
Now, since $\phi(t)=\mathrm{Ad}(\exp(te_{1}))\in{\rm SU}(3)\subset\mathrm{G}_{2}$ [cf. Lemma 2.2], there exist $t_{0}\in\mathbb{R}$ such that
$\varphi=\phi(t_{0})^{*}\widetilde{\varphi}$= $-ze^{1}\wedge(e^{25}+e^{36}+e^{47})+xe^{234}+a(-e^{267}+e^{357}-e^{456})$
$+b(e^{237}-e^{246}+e^{345})-ye^{567},$
satisfies (2.10). On the other hand, the induced metric is given by
$g=\frac{B}{6^{\frac{2}{9}}\mathop{\mathrm{det}}\nolimits(B)^{\frac{1}{9}}},$
hence, by comparing $g$ with (2.7), we obtain the conditions in (2.13). Finally, the expression in (2.11) is obtained by writing $\varphi$ in the orthonormal basis (2.8). ∎
Definition 2.4.
A $\mathrm{G}_{2}$-structure $\varphi$ is called coclosed if $d\psi=0$. Among the class of coclosed class, $\varphi$ is said purely coclosed if $d\varphi\wedge\varphi=0$ and nearly parallel if $d\varphi=4\psi$.
From Lemma (2.3), the following result describes all coclosed invariant $\mathrm{G}_{2}$–metrics on $V^{5,2}$
Theorem 2.1.
The $\mathrm{G}_{2}$-structure (2.11) is coclosed for any $(a,b,x,y,z)\in\mathbb{R}^{5}$ satisfying (2.10). Moreover, $V^{5,2}$ does not carry a ${\rm SO}(5)$-invariant purely coclosed $\mathrm{G}_{2}$-structure and (2.11) is nearly parallel if
$x=a,\quad y=b,\quad a^{2}+b^{2}=\frac{27}{512}\quad\text{and}\quad z=-\frac{9}{32}.$ (2.14)
Proof.
From Lemma 2.3, we get the dual 4-form $\psi=\ast\varphi$
$\psi=$ $-(X^{2356}+X^{2457}+X^{3467})-p_{4}X^{1234}+p_{3}X^{1}\wedge(X^{267}-X^{357}+X^{456})$ (2.15)
$+p_{2}X^{1}\wedge(X^{237}-X^{246}+X^{345})-p_{1}X^{1567}.$
Using the Lie bracket (2.9) and the metric coefficients (2.7), we obtain the exterior derivative for the following $\mathrm{Ad}({\rm SO}(3))$-forms on $\mathfrak{m}$}:
$dX^{1}=$ $\frac{z}{y_{2}y_{3}}\left(X^{25}+X^{36}+X^{47}\right)$ (2.16)
$dX^{234}=$ $-\frac{y_{2}}{z}\left(X^{1237}-X^{1246}+X^{1345}\right),\quad dX^{567}=\frac{y_{3}}{z}\left(X^{1267}-X^{1357}+X^{1456}\right)$
$d(X^{267}$ $-X^{357}+X^{456})=\frac{2y_{3}}{z}\left(X^{1237}-X^{1246}+X^{1345}\right)-\frac{3y_{2}}{z}X^{1567}$
$d(X^{237}$ $-X^{246}+X^{345})=\frac{3y_{3}}{z}X^{1234}-\frac{2y_{2}}{z}\left(X^{1267}-X^{1357}+X^{1456}\right).$
From (2.15) and (2.16), it is easy to check that $d\psi=0$. Now, from (2.11) we obtain
$d\varphi=$ $\frac{2z}{y_{2}y_{3}}\left(X^{2356}+X^{2457}+X^{3467}\right)-\frac{p_{1}y_{2}+2p_{2}y_{3}}{z}\left(X^{1237}-X^{1246}+X^{1345}\right)$ (2.17)
$+\frac{3p_{2}y_{2}}{z}X^{1567}+\frac{3p_{3}y_{3}}{z}X^{1234}-\frac{p_{4}y_{3}+2p_{3}y_{2}}{z}\left(X^{1267}-X^{1357}+X^{1456}\right).$
From a direct computation, we note that $\ast\left(d\varphi\wedge\varphi\right)=12+\frac{3}{y_{2}}+\frac{3}{y_{3}}\neq 0,$ in other words, $\varphi$ given in (2.11) is not purely coclosed. Finally, by imposing the condition $d\varphi=4\psi$, and using (2.15) and (2.17), we obtain (2.14). ∎
Remark 2.5.
The $S^{1}$-family given in (2.14) corresponds to the general construction given in [alexandrov2012]*(4.28).
2.2 Homogeneous Sasakian metrics on $V^{5,2}$
In this section, we study the Sasakian condition for ${\rm SO}(5)$-invariant metrics on $V^{5,2}$. Firstly, we characterise the class $(\eta,\xi,g,\Phi)$ of contact metric structures, cf. Section 1. Secondly, in Theorem 2.2, we provide the constraints in the metric to get a Sasakian structures. And finally, we relate the invariant $\mathrm{G}_{2}$-structures inducing a Sasakian metric.
Lemma 2.6.
The globally defined $1$-form $\eta=X^{1}$ and $\xi=X_{1}$ define a ${\rm SO}(5)$-invariant almost contact structure on $V^{5,2}$. Furthermore, $(\eta,\xi,g,\Phi)$ defines a contact metric structure on $V^{5,2}$, if and only if
$g=4y_{2}y_{3}Q|_{{\mathfrak{m}}_{1}}+y_{2}Q|_{{\mathfrak{m}}_{2}}+y_{3}Q|_{{\mathfrak{m}}_{3}}\quad\text{and}\quad\Phi(X)=[e_{1},X].$ (2.18)
Proof.
First, note that $(V^{5,2},\eta)$ is a contact manifold, since it is clear that $\eta(\xi)=1$ and that $\eta\wedge d\eta^{3}\neq 0$ follows from (2.16). Now, we define an $\mathrm{Ad}({\rm SO}(3))$-invariant contact metric structure $(\eta,\xi,g,\Phi)$, i.e., we define an $\mathrm{Ad}({\rm SO}(3))$-invariant endomorphism $\Phi$ on $\mathfrak{m}$} such that $\Phi^{2}=-I+\eta\otimes\xi$ and $d\eta(X,Y)=2g(X,\Phi(Y))$. From (2.9) we have
$d\eta=\sqrt{\frac{y_{1}}{y_{2}y_{3}}}\left(X^{25}+X^{36}+X^{47}\right),$
Hence, the condition $d\eta(X_{i},X_{i+3})=2g(X_{i},\Phi(X_{i+3}))$ implies that $\Phi$ is given by
$\Phi X_{i}=-\frac{1}{2}\sqrt{\frac{y_{1}}{y_{2}y_{3}}}X_{i+3},\quad\Phi X_{i+3}=\frac{1}{2}\sqrt{\frac{y_{1}}{y_{2}y_{3}}}X_{i},\quad\text{and}\quad\Phi\xi=0.$ (2.19)
Moreover, using the above equalities and the condition $\Phi^{2}=-I+\eta\otimes\xi$, we obtain that $\frac{1}{4}\frac{y_{1}}{y_{2}y_{3}}=-1$, or equivalently $y_{1}=4y_{2}y_{3}$. ∎
The Nijenhuis tensor induced by $\Phi$ is defined by
$N_{\Phi}(X,Y)=[\Phi X,\Phi Y]+\Phi^{2}[X,Y]-\Phi[\Phi X,Y]-\Phi[X,\Phi Y].$ (2.20)
From [Boyer2008]*Theorem 6.5.9, we know that an almost contact structure $(\xi,\eta,\Phi)$ is normal if the Nijenhuis tensor satisfies
$N_{\Phi}=-d\eta\otimes\xi.$ (2.21)
Following [Boyer2008]*Definition 6.5.13, we define:
Definition 2.7.
A normal contact metric structure $\mathcal{S}=(\eta,\xi,g,\Phi)$ on $M^{2n+1}$ is called Sasakian and $(M^{2n+1},\mathcal{S})$ is called a Sasakian manifold.
Theorem 2.2.
The ${\rm SO}(5)$-invariant contact metric structure $(\eta,\xi,g,\Phi)$ given in Lemma 2.6 is a Sasakian structure, if and only if, the metric (2.18) satisfies
$y_{1}=4y_{2}^{2}\quad\text{and}\quad y_{2}=y_{3}.$ (2.22)
Proof.
We show that $(\eta,\xi,g,\Phi)$ is a normal, contact metric structure, i.e., Sasakian. Using the bracket (2.9) and the definition of $\Phi$ in (2.18), we verify that the Nijenhuis tensor (2.20) satisfies (2.21). From Lemma 2.6, we recall that the metric $g$ is contact for $y_{1}=4y_{2}y_{3}$. First, note that $N_{\Phi}(\xi,Y)=\Phi^{2}[\xi,Y]-\Phi[\xi,\Phi Y]$ and now, for $i=1,2,3$ we consider
$\Phi^{2}[\xi,X_{i}]-\Phi[\xi,\Phi X_{i}]$ $=-[\xi,X_{i}]-\Phi[\xi,-X_{i+3}]=-[\xi,X_{i}]+\sqrt{\frac{y_{2}}{y_{1}y_{3}}}\Phi X_{i}$
$=\sqrt{\frac{y_{3}}{y_{1}y_{2}}}X_{i+3}-\sqrt{\frac{y_{2}}{y_{1}y_{3}}}X_{i+3}$
$=\left(\frac{1}{2y_{2}}-\frac{1}{2y_{3}}\right)X_{i+3}$
Since $i_{\xi}d\eta=0$, we conclude that $y_{2}=y_{3}$. Analogously
$\Phi^{2}[\xi,X_{i+3}]-\Phi[\xi,\Phi X_{i+3}]$ $=-[\xi,X_{i+3}]-\Phi[\xi,X_{i}]=\sqrt{\frac{y_{2}}{y_{1}y_{3}}}X_{i}-\sqrt{\frac{y_{3}}{y_{1}y_{2}}}\Phi X_{i+3}$
$=\frac{1}{\sqrt{y_{1}}}X_{i}-\frac{1}{\sqrt{y_{1}}}X_{i}=0.$
For the missing cases, for $X,Y\in{\mathfrak{m}}_{2}\oplus{\mathfrak{m}}_{3}$ the brackets $[X,Y]$, $\Phi Y,Y]$ and $X,\Phi Y]$ are generated by $\xi$ (cf. (2.9)), hence $\Phi[X,Y]=\Phi[\Phi X,Y]=\Phi[X,\Phi Y]=0$ and thus $N_{\Phi}(X,Y)=[\Phi X,\Phi Y]$. So, (2.21) holds by noting that
$N_{\Phi}(X_{i},X_{i+3})=-2X_{1}\quad\text{for}\quad i=2,3,4\quad\text{and}\quad N_{\Phi}(X_{i},X_{j})=0\quad\text{otherwise}.$
According to the notation in Theorem 2.2, the Sasakian metric (2.7) is induced by the following family of $\mathrm{G}_{2}$-structures.
Corollary 2.8.
Consider the Sasakian metric on $V^{5,2}$ as a metric induced by (2.11) and the parameter vector $(z,a,b,x,y)\in\mathbb{R}^{5}$ satisfying the condition (2.10). Then, the $\mathrm{G}_{2}$-structure (2.11) induces a Sasakian metric, if and only if,
$x=a,\quad y=b\quad\text{and}\quad z^{2}=4(a^{2}+b^{2})^{\frac{4}{3}}.$ (2.23)
In this case, the 22-parameter family of coclosed $\mathrm{G}_{2}$-structures is
$\varphi=$ $-X^{1}\wedge(X^{25}+X^{36}+X^{47})+\left(\frac{2}{z}\right)^{\frac{3}{4}}\Big\{a\Big(X^{234}-X^{267}+X^{357}-X^{456}\Big)$
$+b\Big(X^{237}-X^{246}+X^{345}-X^{567}\Big)\Big\},$
$\psi=$ $-\Big(X^{2356}+X^{2457}+X^{3467}\Big)-\left(\frac{2}{z}\right)^{\frac{3}{4}}\Big\{bX^{1}\wedge\Big(X^{234}-X^{267}+X^{357}-X^{456}\Big)$
$-aX^{1}\wedge\Big(X^{237}-X^{246}+X^{345}-X^{567}\Big)\Big\}.$
Proof.
We compare (2.13) and (2.22), thus
$z^{2}=4(a^{2}+by)^{4/3}\quad\text{and}\quad a^{2}+by=b^{2}+ax>0,$
since $a^{2}b^{2}=abxy$, we have $a^{2}(ax+b^{2})=(b^{2}+ax)ax$ and $by(a^{2}+by)=b^{2}(by+a^{2})$. Thus, the result follows.
Remark 2.9.
On the $\mathrm{ad}(\mathfrak{so}(3))$-invariant reductive decomposition ${\mathfrak{m}}={\mathfrak{m}}_{1}\oplus{\mathfrak{m}}_{2}\oplus{\mathfrak{m}}_{3}$ (2.3) the tensors given by
$\xi=\frac{4}{3}e_{1},\quad\Phi=\begin{pmatrix}0&&\\ &&I_{3}\\ &-I_{3}&\end{pmatrix}\quad\text{and}\quad g=\frac{3}{8}\begin{pmatrix}\frac{3}{8}&0\\ 0&I_{6}\end{pmatrix}$,
define the ${\rm SO}(5)$-invariant Sasakian-Einstein structure on $V^{5,2}$, associated with nearly parallel $\mathrm{G}_{2}$-structure of type II (cf.[friedrich1997nearly]*Table 2).
3 $\mathrm{G}_{2}$ and self dual contact instantons on $V^{5,2}$
According to Corollary 2.8, we can consider solutions of (1.2) for those $\mathrm{G}_{2}$-structures with induced Sasakian metric. Additionally, if we endow a complex vector bundle $E\to V^{5,2}$ with a holomorphic Sasakian structure [biswas2010vector], one obtain the notion of Chern connection, as those connections simultaneously compatible with the holomorphic structure and some Hermitian bundle metric. The Hermitian Yang-Mills (HYM) condition is naturally extended to a Sasakian version, namely, let $\omega\colon=d\eta\in\Omega^{1,1}(M)$ be the transversal Kähler form, we say that a connection $A\in\mathcal{A}(E)$ is transverse Hermitian Yang-Mills connection (tHYM) if
$\hat{F}_{A}\colon=\langle F_{A},\omega\rangle=0,\quad\text{and}\quad F^{0,2}_{A}=0.$ (3.1)
Here, the type $(p,q)$ of a $k$-form on $M$ is induced by the condition $\Phi|_{H}^{2}=-1$, where $H$ is the Kähler distribution given by $H=\ker\eta\subset TM=H\oplus\xi$. Additionally, the splitting on the tangent bundle induced by the contact structure yields the splitting $\Omega^{k}(M)=\Omega^{k}(H^{\ast})\oplus\eta\wedge\Omega^{k-1}(H^{\ast})$. Hence, after complexification of the tangent bundle, one can naturally extend this bi-degree $(p,q)$ to the bundle of $E$-valued $k$-forms.
We know that $2$-forms on $M$, and consequently $\Omega^{2}({\mathfrak{g}}_{E})$, decomposes as (cf. [portilla2023, Section. 2.1])
$\Omega^{2}(M)=\underbrace{\Omega^{2}_{1}\oplus\Omega^{2}_{6}\oplus\Omega^{2}_{8}}_{\text{Transversal $2$-forms}}\oplus\Omega^{2}_{V}$ (3.2)
Furthermore, a connection $A\in\mathcal{A}(E)$ is SDCI if its curvature belong to $\Omega^{2}_{8}({\mathfrak{g}}_{E})$, or equivalently, if the curvature $F_{A}\in\Omega^{2}({\mathfrak{g}}_{E})$ is of type $(1,1)$ orthogonal to $\omega$ (cf. [portilla2023, Proposition 2.5]).
Lemma 3.1.
With respect to the Sasakian structure of $V^{5,2}$ given in Theorem 2.2 and the $\mathrm{G}_{2}$-structure in Corollary 2.8 we have that: If AA is a SDCI, then it is a $\mathrm{G}_{2}$-instanton.
Proof.
We know that in local coordinates, a basis for self dual contact 22–forms is given by (cf. [portilla2023, Section 2])
$\begin{array}[]{lccl}w_{1}=X^{23}+X^{56},&w_{2}=X^{26}+X^{35},&w_{3}=X^{24}+X^{57},&w_{4}=X^{27}+X^{45},\\[4.0pt] w_{5}=X^{34}+X^{67},&w_{6}=X^{37}+X^{46},&w_{7}=X^{25}-X^{47},&w_{8}=X^{36}-X^{47}.\end{array}$ (3.3)
For $\mathop{\mathrm{Re}}\Omega_{0}$ and $\mathop{\mathrm{Im}}\Omega_{0}$ given in Lemma 2.1, one verifies that $w_{i}\wedge\real\Omega_{0}=w_{i}\wedge\mathop{\mathrm{Im}}\Omega_{0}=0$ for each $i=1,\cdots,8$ in (3.3). Hence, if $A$ is a SDCI, then its curvature $F_{A}$ is locally generated by $w_{i}$’s in (3.3). Thus, for the $\mathrm{G}_{2}$-structure (cf. Corollary 2.8)
$\varphi=-\eta\wedge d\eta+\left(\frac{2}{z}\right)^{\frac{3}{4}}a\mathop{\mathrm{Re}}(\Omega_{0})+\left(\frac{2}{z}\right)^{\frac{3}{4}}b\mathop{\mathrm{Im}}(\Omega_{0}),$
we have
$\ast(\varphi\wedge F_{A})$ $=-\ast(\eta\wedge d\eta\wedge F_{A})+\left(\frac{2}{z}\right)^{\frac{3}{4}}a\ast(\mathop{\mathrm{Re}}(\Omega_{0})\wedge F_{A})+\left(\frac{2}{z}\right)^{\frac{3}{4}}b\ast(\mathop{\mathrm{Im}}(\Omega_{0})\wedge F_{A})$
$=-\ast(\eta\wedge d\eta\wedge F_{A})=-F_{A}.$
Remark 3.2.
We can generalise Lemma 3.1 in the following sense. Observe that from the decomposition in (3.2), we have the ${\rm SU}(3)$-invariant decomposition
$\Omega^{2}(H)=\Omega^{2}_{6}\oplus\Omega^{2}_{1}\oplus\Omega_{8}^{2},$ (3.4)
where $\Omega^{2}_{8}\simeq\mathfrak{su}(3)$ and $\Omega^{2}_{1}=\Omega^{0}(M)\otimes\omega$. On the other hand, we know that the space of $2$-forms on $H^{\mathbb{C}}=H\otimes_{\mathbb{R}}\mathbb{C}$ decomposes into the $\operatorname{Gl}(3,\mathbb{C})$-invariant submodules
$\Omega^{2}(H^{\mathbb{C}})=\Omega^{2,0}(H^{\mathbb{C}})\oplus\Omega^{0,2}(H^{\mathbb{C}})\oplus\Omega^{1,1}(H^{\mathbb{C}}).$
By Schur’s Lemma, we have (see [falcitelli1994]*Equation (1.2)
$\Omega^{2,0}\oplus\Omega^{0,2}\simeq\Omega^{2}_{6}\otimes_{\mathbb{R}}\mathbb{C}\quad\text{and}\quad\Omega^{1,1}\simeq(\Omega^{2}_{1}\oplus\Omega^{2}_{8})\otimes_{\mathbb{R}}\mathbb{C}.$ (3.5)
On the other hand, we have the $\mathrm{G}_{2}$-invariant decomposition $\Omega^{2}(TM)=\Omega^{2}_{7}\oplus\Omega^{2}_{14},$ where $\Omega^{2}_{14}\simeq{\mathfrak{g}}_{2}$. Since ${\rm SU}(3)\subset\mathrm{G}_{2}$ we have that $\Omega_{8}^{2}\subset\Omega_{14}^{2}$ and $\Omega^{2}_{1}\subset\Omega^{2}_{7}$. It is easy to see that $\Omega^{2}_{V}$ has elements in both $\Omega_{7}^{2}$ and $\Omega_{14}^{2}$. Hence, the same is true for $\Omega^{2}_{6}$.
Lemma 3.3.
Let $E\to V^{5,2}$ be a Sasakian holomorphic vector bundle [Boyer2008, Section 3], the Sasakian structure $(\eta,\xi,g,\Phi)$ is given in Theorem 2.2 and $A\in\mathcal{A}(E)$ the Chern connection. If $A\in\mathcal{A}(E)$ is a $\mathrm{G}_{2}$-instanton, then $A$ satisfies the equation (1.2), i.e. it is a $\lambda$-contact instanton .
Proof.
Consider $\varphi$ given in Corollary 2.8 and set $dz^{k}=X^{k}+\sqrt{-1}X^{k+3}$ for $k=2,3,4$. From Lemma 2.1 and (2.6), we know that $\Omega_{0}=dz^{2}\wedge dz^{3}\wedge dz^{4}\in\Lambda^{3,0}H^{\ast}_{\mathbb{C}}$, where $H:=\ker(\eta)$. Now, if $A\in\mathcal{A}(E)$ is the Chern connection, then $F_{A}$ is of $(1,1)$-type [biswas2010vector, p.552]. By taking into consideration the bi-degree of $\Omega_{0}$, we conclude that $F_{A}\wedge\Omega_{0}=0$ and $F_{A}\wedge\overline{\Omega_{0}}=0$, hence both $F_{A}\wedge\mathop{\mathrm{Im}}{\Omega_{0}}$ and $F_{A}\wedge\mathop{\mathrm{Re}}{\Omega_{0}}$ vanishes, then
$F_{A}\wedge\varphi$ $=F_{A}\wedge\left(\eta\wedge d\eta+\left(\frac{2}{z}\right)^{\frac{3}{4}}a\mathop{\mathrm{Re}}\Omega_{0}+\left(\frac{2}{z}\right)^{\frac{3}{4}}b\mathop{\mathrm{Im}}\Omega_{0}\right)$
$=F_{A}\wedge\eta\wedge d\eta.$
Consequently, solutions to the $\mathrm{G}_{2}$-instanton equation are also solutions to the contact instanton equation. ∎
Consider the Kähler manifold $G_{2}(\mathbb{R}^{5})\cong V^{5,2}/S^{1}$, we fix a holomorphic bundle $(\mathcal{E},\bar{\partial})$ on $G_{5,2}$, the pullback $\mathcal{E}^{\ast}:=\pi^{\ast}\mathcal{E}\to V^{5,2}$ carries a structure of Sasakian holomorphic bundle
[…]
where $\pi\colon V^{5,2}\to G_{2}(\mathbb{R}^{5})$ is a Sasakian circle fibration endowed with the $\mathrm{G}_{2}$-structure (2.11). Finally, in [portilla2023, Lemma 2.4] was proved that SDCI are precisely connections whose curvature is of type $(1,1)$ and orthogonal to $\omega$, (cf. the splitting in (3.2) and Remark 3.2). Thus, from Lemmas 3.1 and 3.3, we obtain the following result:
Corollary 3.4.
The vector bundle $\mathcal{E}^{\ast}$ is a holomorphic Sasakian bundle, and a Chern connection $A$ on $\mathcal{E}$ is HYM if, and only if, $\pi^{\ast}A$ is a $\mathrm{G}_{2}$-instanton on $\mathcal{E}^{*}$.
We recall that in Lemma 3.1, Lemma 3.3 and Corollary 3.4 the key argument follow just in presence of a SU(n)(n)-structure. These have analogous results in [portilla2023, Proposition 3.11] and [calvo2020gauge, Lemma 21], the difference in our case is that we do not assume $M$ is a contact Calabi-Yau but only endowed with a transversal ${\rm SU}(3)$-structure.
4 Homogeneous Principal Bundles on $V^{5,2}$
First, we recall some notation and results about homogeneous principal bundles, following [cap2009]*Section 1.4. Let $H\subset G$ be a closed subgroup of the Lie group $G$, we denote by $M=G/H$ the homogeneous space, let $L\colon G\times M\to M$ be the canonical left action and $p\colon G\to M$ an HH–principal bundle given by the canonical projection.
Definition 4.1.
A homogeneous principal bundle over $M$ is a locally trivial principal bundle $\pi\colon P\to M$ together with a left GG–action $\widetilde{L}\colon G\times P\to P$, which lifts the action $L$ on $M$, i.e., $\pi(\widetilde{L}({\textsl{g}},y))=L({\textsl{g}},\pi(y))$, and for each ${\textsl{g}}\in G$, the bundle map $\widetilde{L}_{\textsl{g}}\colon P\to P$ is a homomorphism of principal bundles.
Homogeneous principal bundles are characterised in the following way:
Lemma 4.2.
[cap2009, Lemma 1.4.5] Let $G$ and $K$ be Lie groups and $H\subset G$ a closed Lie subgroup. Let $P\to G/H$ be a homogeneous $K$-principal bundle, then there exists a smooth homomorphism $\phi\colon H\to K$ such that $P\cong G\times_{\phi}K$. The action of $H$ on $K$ is given by $h\cdot k=\phi(h)\cdot k$ for each $h\in H$ and $k\in K$. The bundles corresponding to two homomorphisms $\phi,\widetilde{\phi}\colon H\to K$ are isomorphic if and only if there exists $k\in K$ such that $k\cdot\phi=\widetilde{\phi}\cdot k$
For a homogeneous $K$-principal bundle $P\to G/H$, a connection $A\in\Omega^{1}(P,{\mathfrak{k}})$ is called invariant if $\widetilde{L}_{\textsl{g}}^{*}A=A$ for each ${\textsl{g}}\in G$.
Lemma 4.3.
[cap2009, Theorem 1.4.5] Let $P=G\times_{\phi}K\to G/H$ be the homogeneous $K$-principal bundle induced by the homomorphism $\phi:H\to K$. Then, invariant principal connections on $P$ are in bijective correspondence with linear maps $\alpha\colon{\mathfrak{g}}\to{\mathfrak{k}}$ such that
  • (i)
    $\alpha|_{{\mathfrak{h}}}=\phi_{\ast}\colon{\mathfrak{h}}\to{\mathfrak{k}}$, the derivative of $\phi$.
  • (ii)
    $\alpha\circ\mathrm{Ad}(h)=\mathrm{Ad}(\phi(h))\circ\alpha$ for all $h\in H$.
4.1 Homogeneous ${\rm SO}(3)$-principal bundles
Consider $H=K={\rm SO}(3)$ and $G={\rm SO}(5)$. Since ${\rm SO}(3)$ is simple, up to conjugation there are two smooth homomorphisms $\phi_{i}\colon{\rm SO}(3)\to{\rm SO}(3),$ $i=0,1$. Namely, $\phi_{0}$ the trivial homomorphism and $\phi_{1}=1_{{\rm SO}(3)}$ the identity homomorphism, according to Lemma 4.2, there exist precisely two [^3] Up to isomorphism of homogeneous principal bundles homogeneous ${\rm SO}(3)$-principal bundles $\pi_{i}\colon P_{i}=G\times_{\phi_{i}}{\rm SO}(3)\to V^{5,2}$, $i=0,1$.
Lemma 4.4.
There is a unique invariant ${\rm SO}(3)$–connection on $P_{0}=V^{5,2}\times{\rm SO}(3)$, given by
$\alpha=e^{1}\otimes(b_{8}e_{8}+b_{9}e_{9}+b_{10}e_{10}).$ (4.1)
Furthermore, the curvature $F_{\alpha}$ of the connection $\alpha$ belongs to $\Omega^{2}_{1}(\mathfrak{so}(3))$ cf. (3.4) and Remark 3.2.
Proof.
Consider an arbitrary linear map $\alpha:\mathfrak{so}(5)\rightarrow\mathfrak{so}(3)$, in the basis (2.2) we write
$\alpha=a_{i}e^{i}\otimes e_{8}+b_{j}e^{j}\otimes e_{9}+c_{k}e^{k}\otimes e_{10}.$
From Lemma 4.3, $\alpha$ is an invariant connection, if and only if, it satisfies $\alpha|_{\mathfrak{so}(3)}=0$ and $\alpha\circ\mathrm{Ad}({\textsl{g}})=\alpha$, for all ${\textsl{g}}\in{\rm SO}(3)$. The condition $\alpha|_{\mathfrak{so}(3)}=0$ implies that $\alpha$ is of the form $\alpha=a_{i}e^{i}\otimes e_{8}+b_{j}e^{j}\otimes e_{9}+c_{k}e^{k}\otimes e_{10}$ for $i,j,k\in\{1,\dots,7\}$. Now, we apply the second condition to ${\textsl{g}}_{k}=\begin{pmatrix}I_{2\times 2}&\\ 0&h_{k}\\ \end{pmatrix}\in{\rm SO}(5),$ we define some values of $h_{k}$,
$h_{0}=\begin{pmatrix}0&1&0\\ 1&0&0\\ 0&0&-1\end{pmatrix},h_{1}=\begin{pmatrix}0&1&0\\ -1&0&0\\ 0&0&1\end{pmatrix},h_{2}=\begin{pmatrix}0&0&1\\ 0&-1&0\\ 1&0&0\end{pmatrix}\quad\text{and}\quad h_{3}=\begin{pmatrix}0&0&1\\ 0&1&0\\ -1&0&0\end{pmatrix}$ (4.2)
First, considering $h_{0}$ and $h_{1}$, we obtain that $\alpha(e_{i})=a_{i}e_{8}+b_{i}e_{9}+c_{i}e_{10}$ and, for $i\neq 3$ and $i\neq 6$
$\alpha(\mathrm{Ad}({\textsl{g}}_{0})(e_{i}))=\alpha({\textsl{g}}_{0}e_{i}{\textsl{g}}_{0}^{-1})$ $=a_{i}e_{8}+b_{i}e_{10}+c_{i}e_{9}$
$\alpha(\mathrm{Ad}({\textsl{g}}_{1})(e_{i}))=\alpha({\textsl{g}}_{1}e_{i}{\textsl{g}}_{1}^{-1})$ $=-a_{i}e_{8}-b_{i}e_{10}-c_{i}e_{9}$
it follows that $a_{i}=b_{i}=c_{i}=0$ for $i\in\{2,4,5,7\}$. Now, by considering $h_{2}$ and $h_{3},$ we obtain for $i=3,6$
$\begin{array}[]{cl}\alpha(\mathrm{Ad}({\textsl{g}}_{0})(e_{i}))=&-\alpha(e_{i})=a_{i}e_{8}+b_{i}e_{10}+c_{i}e_{9}\\[5.0pt] \alpha(\mathrm{Ad}({\textsl{g}}_{1})(e_{i}))=&\alpha(e_{i})=-a_{i}e_{8}-b_{i}e_{10}-c_{i}e_{9}.\end{array}$
Consequently, an invariant connection $\alpha$ is of the form
$\alpha=e^{1}\otimes(b_{8}e_{8}+b_{9}e_{9}+b_{10}e_{10}).$ (4.3)
The curvature of $\alpha$ is $F_{\alpha}=y_{2}^{-1}(X^{25}+X^{36}+X^{47})\otimes(b_{8}e_{8}+b_{9}e_{9}+b_{10}e_{10})$. The second statement in the Lemma follows from (2.6) since
$F_{\alpha}=2\omega\otimes(b_{8}e_{8}+b_{9}e_{9}+b_{10}e_{10}).$ (4.4)
Lemma 4.5 (Homogeneous ${\rm U}(1)$-connections).
There is a unique invariant ${\rm U}(1)$-connections on $P=V^{5,2}\times{\rm U}(1)$, given by
$\beta=a_{1}e^{1}\otimes e_{1}$ (4.5)
Furthermore, the curvature $F_{\alpha}$ of the connection $\alpha$ belongs to $\Omega^{2}_{1}$ cf. (3.4).
Proof.
Similar to Lemma 4.4, since ${\rm SO}(3)$ is simple, all non-trivial homomorphism $\phi\colon{\rm SO}(3)\rightarrow K$ is either trivial or injective for all Lie group $K$. In particular for $K={\rm U}(1)$, from Lemma 4.2, there exists a unique [^4] Up to isomorphism of principal bundles homogeneous ${\rm U}(1)$-bundle $P$ corresponding to $\phi$, namely, the trivial bundle $P=V^{5,2}\times{\rm U}(1)$. Any invariant connection on $P$ is given by a linear map $\beta:\mathfrak{so}(5)\rightarrow{\mathfrak{u}}(1)$ satisfying
$(i)\quad\beta|_{\mathfrak{so}(3)}=0=\phi_{\ast}\colon\mathfrak{so}(3)\to\mathbb{R}\quad\text{and}\quad(ii)\quad\beta\circ\mathrm{Ad}(a)=\beta,\text{ for all }a\in{\rm SO}(3).$
By considering $e_{1}$ the generator of the Abelian Lie algebra ${\mathfrak{u}}(1)=\mathbb{R}$, the unique invariant connection is given by
$\alpha=a_{i}e^{i}\otimes e_{1}\quad\text{for}\quad i=1,\dots,10.$
From item $(i)$, since $\alpha|_{\mathfrak{so}(3)}=0$, $\alpha$ has to be of the form $\alpha=a_{i}e^{i}\otimes e_{1}$, $i=1,\cdots,7.$ On the other hand, considering item $(ii)$ for ${\textsl{g}}_{0},{\textsl{g}}_{1}$ given in (4.2), we use the same argument used in the proof of Lemma 4.4 to obtain $\alpha(e_{i})=a_{i}e_{1}$ and for $i=2,\cdots,7$
$\alpha(\mathrm{Ad}({\textsl{g}}_{0})(e_{i}))=\alpha({\textsl{g}}_{0}e_{i}{\textsl{g}}_{0}^{-1}))$ $=-a_{i}e_{1}$
$\alpha(\mathrm{Ad}({\textsl{g}}_{1})(e_{i}))=\alpha({\textsl{g}}_{1}e_{i}{\textsl{g}}_{1}^{-1}))$ $=a_{i}e_{1}.$
Therefore, $\alpha$ is characterised by
$\beta=a_{1}e^{1}\otimes e_{1}$ (4.6)
and its curvature $2$-form is
$F_{\beta}=2(X^{25}+X^{36}+X^{47})\otimes e_{1}.$ (4.7)
Note that $F_{\beta}\in\Omega^{2}_{1}=\omega\otimes\Omega^{0}(M)$ (4.14). ∎
Now, we consider the ${\rm SO}(3)$-principal bundle $P_{1}={\rm SO}(5)\times_{\phi_{1}}{\rm SO}(3)$ induced by the homomorphism identity $\phi_{1}\colon{\rm SO}(3)\to{\rm SO}(3)$. The projection of $P_{1}$ is given by
$\pi_{1}\colon[(u,s)]\in P_{1}\mapsto u\cdot{\rm SO}(3)\in V^{5,2},$ (4.8)
and the class is given by $(u,s)]=\{(u\cdot r,r^{-1}s)\;|\;r\in{\rm SO}(3)\}.$ There exist a left ${\rm SO}(5)$-action on $P_{1}$ such that for each $v\in{\rm SO}(5)$ the bundle map $L_{v}\colon P_{1}\to P_{1}$ is $\pi_{1}$-equivariant, i.e.,
$\pi_{1}(L_{v}[(u,s)])=\pi_{1}([(vu,s)])=vu\cdot{\rm SO}(3)=L_{v}(\pi_{1}[(u,s)]).$
According to (2.2), we fix the basis $\{e_{8},e_{9},e_{10}\}$ for $\mathfrak{so}(3)\subset\mathfrak{so}(5)$ such that $e_{i},e_{j}]=\varepsilon_{ijk}e_{k}$, where $\varepsilon_{ijk}$ is the sign of the permutation (ijk)(i\;j\;k).
Lemma 4.6.
There is a unique invariant ${\rm SO}(3)$–connection on $P_{1}=V^{5,2}\times_{\phi_{1}}{\rm SO}(3)$ given by
$\alpha=e^{8}\otimes e_{8}+e^{9}\otimes e_{9}+e^{10}\otimes e_{10}.$ (4.9)
Proof.
Consider $\alpha:\mathfrak{so}(5)\rightarrow\mathfrak{so}(3)$ in the basis (2.2) given by $\alpha=a_{i}e^{i}\otimes e_{8}+b_{j}e^{j}\otimes e_{9}+c_{k}e^{k}\otimes e_{10}$. From Lemma 4.3 (i), we have that $\alpha=\alpha_{0}+e^{8}\otimes e_{8}+e^{9}\otimes e_{9}+e^{10}\otimes e_{10},$ where $\alpha_{0}=a_{i}e^{i}\otimes e_{8}+b_{j}e^{j}\otimes e_{9}+c_{k}e^{k}\otimes e_{10}$, with $i,j,k\in\{1,\dots,7\}$, we are going to conclude that $\alpha_{0}=0$. From Lemma 4.3 (ii), for any ${\textsl{g}}\in{\rm SO}(3)$ and $X\in\mathfrak{so}(5)$, we have ${\textsl{g}}\alpha(X){\textsl{g}}^{-1}=\alpha({\textsl{g}}X{\textsl{g}}^{-1})$. Taking $X=e_{i}$ for each $i\in\{1,\dots,7\}$, we denote
$\alpha(\mathrm{Ad}({\textsl{g}})(e_{i}))=\alpha({\textsl{g}}e_{i}{\textsl{g}}^{-1})$ (4.10)
and
$\mathrm{Ad}({\textsl{g}})(\alpha(e_{i}))=a_{i}{\textsl{g}}e_{8}{\textsl{g}}^{-1}+b_{i}{\textsl{g}}e_{9}{\textsl{g}}^{-1}+c_{i}{\textsl{g}}e_{10}{\textsl{g}}^{-1}.$ (4.11)
Now, consider ${\textsl{g}}_{k}$ and $h_{k}$ as in (4.2), equations (4.10) and (4.11) become
$\alpha(\mathrm{Ad}({\textsl{g}}_{0})(e_{2}))=\alpha({\textsl{g}}_{0}e_{2}{\textsl{g}}_{0}^{-1})=\alpha(e_{3})=a_{3}e_{8}+b_{3}e_{9}+c_{3}e_{10},$
and
$\mathrm{Ad}({\textsl{g}}_{0})(\alpha(e_{2}))=-a_{2}e_{8}+b_{2}e_{10}-c_{2}e_{9}.$
These imply $a_{2}=a_{3},-c_{2}=b_{3},c_{3}=b_{2}$. Analogously, for $h_{1}$, we have
$\alpha(\mathrm{Ad}({\textsl{g}}_{1})(e_{2}))=\alpha(e_{3})=-a_{3}e_{8}-b_{3}e_{9}-c_{3}e_{10}$
and
$\mathrm{Ad}({\textsl{g}}_{1})(\alpha(e_{2}))=a_{2}e_{8}+b_{2}e_{10}-c_{2}e_{9},$
consequently, $a_{2}=a_{3},c_{2}=b_{3},-b_{2}=c_{3}$, hence $c_{2}=b_{3}=c_{3}=b_{2}=0$. Similarly:
  • From $\mathrm{Ad}({\textsl{g}}_{0})(\alpha(e_{4}))=\alpha(\mathrm{Ad}({\textsl{g}}_{0})(e_{4}))$, we obtain $c_{4}=b_{4}=0$,
  • From $\mathrm{Ad}({\textsl{g}}_{1})(\alpha(e_{7}))=\alpha(\mathrm{Ad}({\textsl{g}}_{1})(e_{7}))$, we obtain $c_{7}=b_{7}=0$.
  • From $\mathrm{Ad}({\textsl{g}}_{0})(\alpha(e_{5}))=\alpha(\mathrm{Ad}({\textsl{g}}_{0})(e_{5}))$ and $\mathrm{Ad}({\textsl{g}}_{1})(\alpha(e_{5}))=\alpha(\mathrm{Ad}({\textsl{g}}_{1})(e_{5}))$, we obtain $c_{6}=b_{5}=c_{5}=b_{6}=0$.
Finally, applying the same argument for $h_{2}$ and $h_{3},$ we conclude that $a_{i}=0$ for $i=1,\cdots,7$, consequently $\alpha_{0}=0$, then $\alpha=e^{8}\otimes e_{8}+e^{9}\otimes e_{9}+e^{10}\otimes e_{10}$ and (4.9) defines a unique ${\rm SO}(5)$-invariant connection $\alpha\in\Omega^{1}(P,\mathfrak{so}(3))$. ∎
In the basis (2.8), the curvature of (4.9) is given by
$F_{\alpha}=\frac{1}{y_{2}}\left((X^{23}+X^{56})\otimes e_{8}+(X^{24}+X^{57})\otimes e_{9}+(X^{34}+X^{67})\otimes e_{10}\right).$ (4.12)
Notice that $F_{\alpha}:{\mathfrak{m}}\times{\mathfrak{m}}\rightarrow\mathfrak{so}(3)$ is a ${\rm SO}(3)$-equivariant map, i.e., $F_{\alpha}\circ\mathrm{Ad}({\textsl{g}})=\mathrm{Ad}({\textsl{g}})\circ F_{\alpha}$ for any ${\textsl{g}}\in{\rm SO}(3)$, thus (4.12) corresponds with the ${\rm SO}(5)$-invariant curvature $2$-form of the induced by the ${\rm SO}(5)$-invariant connection $\alpha\in\Omega^{1}(P,\mathfrak{so}(3))$ (cf. [cap2009, Proposition 1.4.6]).
4.2 Homogeneous $\mathrm{G}_{2}$ and self dual contact instantons (SDCI)
For the ${\rm SO}(5)$-invariant contact form $\eta=X^{1}$ on $V^{5,2}$ and the ${\rm SO}(5)$-invariant metric given by (2.7), we consider the equation
$\ast_{g}(F_{\beta}\wedge\eta\wedge\omega)=\pm F_{\beta},\quad\text{where}\quad\omega=\frac{1}{2}d\eta=(X^{25}+X^{36}+X^{47}).$ (4.13)
From Lemma 4.6, we obtain the following
Corollary 4.7.
Consider $V^{5,2}$ with the Sasakian structure given in Theorem 2.2. Then, the connection (4.9) is an invariant SDCI on the homogeneous ${\rm SO}(3)$–principal bundle $P_{1}\to V^{5,2}$.
Proof.
For any Sasakian manifold, the bundle of 22–forms decomposes into [portilla2023, Section 2.1]
$\Omega^{2}=\Omega^{2}_{1}\oplus\Omega^{2}_{6}\oplus\Omega^{2}_{8}\oplus\Omega^{2}_{V},$ (4.14)
where (3.3) are elements of $\Omega^{2}_{8}=span\{w_{i}\}_{i=1}^{8}$ solving $\ast(\eta\wedge\omega\wedge w_{i})=w_{i}$. Comparing (4.12) with the values of $w_{i}$ in (3.3), we note that $F_{\alpha}$ in (4.12) is given by
$F_{\alpha}:=\frac{1}{y_{2}}(w_{1}\otimes e_{8}+w_{3}\otimes e_{9}+w_{5}\otimes e_{10}).$
The curvature tensor $F_{\alpha}$ satisfies $\ast(F_{\alpha}\wedge\eta\wedge\omega)=F_{\alpha}$. Hence, for any $y_{2}\in\mathbb{R}^{+}$ satisfying (2.22), the connection (4.9) is an ${\rm SO}(5)$-invariant SDCI. ∎
Corollary 4.8.
Consider $V^{5,2}$ endowed with the Sasakian structure given in Corollary 2.8. There exist a unique invariant ASDCI on the trivial homogeneous ${\rm SO}(3)$–principal bundle $P_{0}\to V^{5,2}$ (cf. Lemma 4.4).
Proof.
From Lemma 4.4, we only should note that the curvature (4.4) satisfy $\ast_{g}(F_{\beta}\wedge\eta\wedge\omega)=F_{\beta},$ if and only if $y_{1}=\frac{9}{2}$. ∎
A result analogous to Corollary 4.8 follows for $P\to V^{5,2}$ in Remark 4.5, since the curvature tensor in (4.7) satisfy $\ast_{g}(F_{\beta}\wedge\eta\wedge\omega)=F_{\beta},$ if and only if $y_{1}=\frac{9}{2}$.
Proposition 4.9.
Consider the connection (4.9) on $V^{5,2}$, with respect to the $\mathrm{G}_{2}$–structure given in Corollary (2.8). Then, we have
  • The ${\rm SO}(3)$–connection in (4.9) on the homogeneous principal bundle $P_{1}\to V^{5,2}$ [cf. Lemma 4.6] satisfies $F_{\alpha}\wedge\psi=0$ for any $y_{1}>0$, i.e., there exist a unique ${\rm SO}(5)$–invariant connection which is simultaneously $\mathrm{G}_{2}$ and SDCI, given by (4.12).
  • The ${\rm SO}(3)$–connection (4.1) on the trivial bundle $P_{0}\to V^{5,2}$ (cf. Lemma 4.4), whose curvature is given in (4.4) satisfies $F_{\alpha}\wedge\psi=0$, for $\psi$ given in (2.15), if and only if (4.1) is flat.
4.3 Yang–Mills connections
Now, we analyse the Yang–Mills condition for instanton (4.9), for a detailed study on stability of Yang–Mills connections we refer to [bourguignon1981stability]. On the space of connections $\mathcal{A}$, the global inner product $(\alpha,\beta)=\int_{M}\langle\alpha\wedge\ast\beta\rangle$ allow us to define the Yang–Mills functional as the $L^{2}$-norm of the curvature:
$\mathcal{YM}\colon\nabla\in\mathcal{A}\to\|{\rm F}_{\nabla}\|^{2}_{L^{2}}=\int_{M}\operatorname{Tr}({\rm F}_{\nabla}\wedge\ast{\rm F}_{\nabla})=\int_{M}\|{\rm F}_{\nabla}\|^{2}\in\mathbb{R}.$ (4.15)
A critical point of (4.15) is called a Yang–Mills (YM) connection, i.e. $\nabla$ satisfies the so-called Yang–Mills equation $d^{\ast}_{\nabla}F_{\nabla}=0$. Taking the exterior derivative $d_{\nabla}$ in (1.2) and using the Bianchi identity, one notes that a solution of the instanton equation (1.2) satisfy the YM-equation with torsion:
$d_{\nabla}(\ast F_{\nabla})=d\sigma\wedge F_{\nabla}.$
Note that neither $\sigma=\eta\wedge d\eta$ nor $\sigma=\varphi$ are closed, consequently, the torsion term in the above equation is not automatically zero, however, we have that SDCI and ASDCI are Yang–Mills connections [luis2024Weitzenbock, Section 5]. With respect to the $\mathrm{G}_{2}$-structure (2.23), we have the following:
Theorem 4.1.
Let $\alpha$ be the unique $\mathrm{G}_{2}$-instanton (4.9) on the homogeneous ${\rm SO}(3)$-principal bundle $P_{1}$. Then, $\alpha$ is a critical point for the Yang–Mills energy for any coclosed $\mathrm{G}_{2}$-structure (2.23).
Proof.
Using (2.17) and the fact that $\alpha$ is a $\mathrm{G}_{2}$-instanton, we have that the curvature (4.12) satisfies $d_{\alpha}^{*}F_{\alpha}=0$, if and only if
$2a^{3}-2a^{2}x-3ab^{2}+2aby+bxy=0,\quad\text{and}\quad 2b^{3}-2b^{2}y-3a^{2}b+2abx+axy=0,$
we see that both equations hold for (2.23), i.e. $x=a$ and $y=b$. Hence, we obtain the result. ∎
5 Infinitesimal deformations and rigidity of $\mathrm{G}_{2}$-instantons
In this section, we extend the spinorial characterisation of infinitesimal deformations of $\mathrm{G}_{2}$–instantons obtained in [singhal2022] to the setting of coclosed $\mathrm{G}_{2}$–structures. As an application, we show that for certain values of the parameter $y_{2}$ appearing in (4.12), the corresponding $\mathrm{G}_{2}$–instanton is rigid. Throughout this section, $(M^{7},\varphi)$ denotes a 77–dimensional manifold endowed with a coclosed $\mathrm{G}_{2}$–structure, and $\zeta$ the unit spinor associated with $\varphi$.
Let us recall, consider $\zeta$ the unit spinor associated to $\varphi$ and the induced spinor representation $\Delta_{7}\simeq\langle\zeta\rangle\oplus TM$, the Clifford product of a vector field $Y$ and a spinor $(f,Z)$ is (see [karigiannis2006some])
$Y\cdot(f,Z)=(-\langle Y,Z\rangle,fY+Y\times Z),$ (5.1)
where $(f,0)=f\zeta$ and $Y\times Z=(\varphi(Y,Z,\cdot))^{\sharp}$. The Clifford product (5.1) is extended to a $k$-form $\varepsilon$ by
$\varepsilon\cdot(f,Y)=\frac{1}{k!}\varepsilon_{i_{1},\dots i_{k}}e_{i_{1}}\cdot(e_{i_{2}}\cdot(\dots\cdot(e_{i_{k}}\cdot(f,Z)\dots))$ (5.2)
Hence, we have the characterisation of the instanton condition (1.3) in the $\mathrm{G}_{2}$ case in terms of spinors:
Definition 5.1.
Let $P\to M$ be a principal bundle. A connection $A$ on a principal bundle $P$ is called a $\mathrm{G}_{2}$–instanton if its curvature satisfies $F_{A}\cdot\zeta=0,$ where $\cdot$ denotes Clifford multiplication (5.2).
5.1 Infinitesimal deformation of $\mathrm{G}_{2}$-instantons
Definition 5.2.
Let $A$ be a $\mathrm{G}_{2}$-instanton on a principal bundle $P\to M$. An infinitesimal deformation of $A$ is a one–form $\varepsilon\in\Omega^{1}(M,\mathrm{Ad}(P))$ satisfying
$(d_{A}\varepsilon)\cdot\zeta=0,\qquad d_{A}^{\ast}\varepsilon=0.$ (5.3)
Let $\nabla^{g}$ be the Levi–Civita connection of $g$ associated to $\varphi$, we consider the one-parameter family of connections
$\nabla^{t}=\nabla^{g}+\frac{t}{2}\,g^{-1}H_{\varphi}\quad\text{for}\quad t\in\mathbb{R},$ (5.4)
and
$H_{\varphi}=\frac{\tau_{0}}{6}\,\varphi-\tau_{3}.$
For $t=1$, it turns out that $\nabla^{1}$ is the unique connection with parallel skew-symmetric torsion $H_{\varphi}$ that makes $\varphi$ parallel [friedrich2002]. The connection $\nabla^{t}$ induces a connection on the spinor bundle $\not${S} such that:
[…]
Definition 5.3.
With respect to the Clifford multiplication $\cdot\colon TM\times\not$$\to\not{S}$, we define the Dirac operator
$D^{t}\mu=\sum_{i=1}^{7}e_{i}\cdot\nabla^{t}_{e_{i}}\mu,$ (5.6)
[…]
Lemma 5.4.
Let $\zeta$ be the spinor associated with $\varphi$, then it holds
$(X\lrcorner\varphi)\cdot\zeta=3X\cdot\zeta,\qquad(X\lrcorner\gamma)\cdot\zeta=2h(X)\cdot\zeta\quad\text{and}\quad(X\lrcorner\ast\gamma)\cdot\zeta=-2h(X)\cdot\zeta,$
where $X\in\mathscr{X}(M)$, $\gamma\in\Omega^{3}_{27}$ and $h$ is the traceless symmetric tensor defining $\gamma$ by
$\gamma_{ijk}=h_{ip}\gamma_{pjk}+h_{jp}\gamma_{ipk}+h_{kp}\gamma_{ijp}.$
Proof.
Now, using (5.2), we get
[…]
$=3g_{im}\,(0,e^{m})=3\,(0,e_{i})=3e_{i}(0,1)$
$=3e_{i}\cdot\zeta$
Here, we used the identity $\varphi_{ijk}\varphi_{mnk}=g_{im}g_{jn}-g_{in}g_{jm}+\psi_{ijmn}$ [karigiannis2009flows]*Lemma A.8. Assume $\gamma\in\Lambda^{3}_{27}$, we compute
$(X\;\lrcorner\;\gamma)\cdot\zeta$ $=X_{i}\,(e_{i}\;\lrcorner\;\gamma)\cdot\zeta=\frac{1}{2}\,X_{i}\,\gamma_{ijk}\,e^{j}\cdot e^{k}\cdot\zeta$
$=\frac{1}{2}\,X_{i}\bigl(h_{i}{}^{p}\,\varphi_{pjk}+h_{j}{}^{p}\,\varphi_{ipk}+h_{k}{}^{p}\,\varphi_{ijp}\bigr)e^{j}\cdot e^{k}\cdot\zeta.$
$=\frac{1}{2}\,X_{i}\bigl(h_{i}{}^{p}\,\varphi_{pjk}+2h_{j}{}^{p}\,\varphi_{ipk}\bigr)e^{j}\cdot e^{k}\cdot(1,0)$
$=\frac{1}{2}\,X_{i}\bigl(h_{i}{}^{p}\,\varphi_{pjk}+2h_{j}{}^{p}\,\varphi_{ipk}\bigr)(0,\,e^{j}\times e^{k})$
$=\frac{1}{2}\,X_{i}\bigl(h_{i}{}^{p}\,\varphi_{pjk}+2h_{j}{}^{p}\,\varphi_{ipk}\bigr)(0,\,\varphi^{jk}{}_{m}e^{m})$
$=\frac{1}{2}\,X_{i}\,h_{i}{}^{p}\,(6\,g_{pm})\,(0,e^{m})+X_{i}\,h_{j}{}^{p}\bigl(g_{im}g^{j}{}_{p}-g_{i}{}^{j}g_{pm}+\psi_{ipm}{}^{j}\bigr)(0,e^{m})$
$=3\,h(X)\cdot\zeta-h(X)\cdot\zeta.$
$=2\,h(X)\cdot\zeta$
The next result characterises the infinitesimal deformations of a $\mathrm{G}_{2}$-instanton in terms of the operator (5.7).
Proposition 5.5.
Let $\varepsilon\in\Omega^{1}(M,\mathrm{Ad}(P))$ and $D^{t,A}$ the Dirac operator (5.7) with $A$ a $\mathrm{G}_{2}$-instanton. Then $\varepsilon$ defines an infinitesimal deformation of AA (cf. Definition 5.2), if and only if
$D^{t,A}(\varepsilon\cdot\zeta)=\frac{(-5+t)}{8}\,\tau_{0}\,\varepsilon\cdot\zeta+(3t-1)\,(\varepsilon\;\lrcorner\;\tau_{27})\cdot\zeta.$
Proof.
Let $\{e_{i}\}_{i=1}^{7}$ be a local orthonormal frame on $M$. By definition of the twisted Dirac operator associated with the product connection $\nabla^{0,A}$, a direct computation gives
$D^{0,A}(\varepsilon\cdot\zeta)$ $=\sum e_{i}\cdot\nabla^{0,A}_{e_{i}}(\varepsilon\cdot\zeta)=(d_{A}\varepsilon+d_{A}^{\ast}\varepsilon)\cdot\zeta+\sum e_{i}\cdot\varepsilon\cdot\nabla^{0,A}_{e_{i}}(\zeta)$
$=(d_{A}\varepsilon+d_{A}^{*}\varepsilon)\cdot\zeta-\sum_{i}\left(\frac{\tau_{0}}{24}e_{i}\cdot\varepsilon\cdot(e_{i}\lrcorner\varphi)\cdot\zeta-\frac{1}{4}e_{i}\cdot\varepsilon\cdot(e_{i}\lrcorner\tau_{3})\cdot\zeta\right).$
Here, we used that $X\cdot Y+Y\cdot X=-2g(X,Y)1$ and $\nabla^{1}\zeta=0$ for the connection (5.5). Now, using Lemma 5.4 we get
$D^{0,A}(\varepsilon\cdot\zeta)$ $=(d_{A}\varepsilon+d_{A}^{*}\varepsilon)\cdot\zeta-\frac{5\tau_{0}}{8}\,\varepsilon\cdot\zeta-(\varepsilon\lrcorner\tau_{27})\cdot\zeta.$
On the other hand, for $H_{\varphi}=\frac{\tau_{0}}{6}\,\varphi-\tau_{3}$ we have
[…]
$=\frac{1}{2}\sum_{i,j,k}H_{ijk}\,\varepsilon_{\ell}\cdot e^{i}\cdot e^{\ell}\cdot e^{j}\cdot e^{k}\cdot\zeta+\sum H_{i\ell k}\varepsilon_{\ell}e^{i}\cdot e^{k}\cdot\zeta-2\,(\varepsilon\;\lrcorner\;H)\cdot\zeta$
$=-\frac{1}{2}\sum H_{ijk}\varepsilon_{\ell}\cdot e^{\ell}\cdot e^{i}\cdot e^{j}\cdot e^{k}\cdot\zeta-\sum H_{\ell jk}\,\varepsilon_{\ell}e^{j}\cdot e^{k}\cdot\zeta-4\,(\varepsilon\;\lrcorner\;H)\cdot\zeta.$
$=-3\,\varepsilon\cdot H\cdot\zeta-6\,(\varepsilon\;\lrcorner\;H)\cdot\zeta=-\frac{1}{2}\,\tau_{0}\,\varepsilon\cdot\varphi\cdot\zeta-6\,(\varepsilon\;\lrcorner\;H)\cdot\zeta.$
$=\frac{7}{2}\,\tau_{0}\,\varepsilon\cdot\zeta-\tau_{0}\,(\varepsilon\;\lrcorner\;\varphi)\cdot\zeta+6\,(\varepsilon\;\lrcorner\;\tau_{3})\cdot\zeta.$
$=\frac{1}{2}\,\tau_{0}\,\varepsilon\cdot\zeta+12\,(\varepsilon\;\lrcorner\;\tau_{27})\cdot\zeta.$
Consequently, from (5.7) we obtain
$D^{t,A}(\varepsilon\cdot\zeta)=(d_{A}\varepsilon+d_{A}^{*}\varepsilon)\cdot\zeta-\frac{5}{8}\,\tau_{0}\,\varepsilon\cdot\zeta-(\varepsilon\;\lrcorner\;\tau_{27})\cdot\zeta+\frac{t}{8}\,\tau_{0}\,\varepsilon\cdot\zeta+3t\,(\varepsilon\;\lrcorner\;\tau_{27})\cdot\zeta.$
Assume that $\varepsilon$ satisfies the linearised instanton equation together with the Coulomb gauge condition, i.e., $d_{A}\varepsilon=0$ and $d_{A}^{*}\varepsilon=0$, then
$D^{t,A}(\varepsilon\cdot\zeta)=\frac{-5+t}{8}\,\tau_{0}\,\varepsilon\cdot\zeta+(3t-1)\,(\varepsilon\;\lrcorner\;\tau_{27})\cdot\zeta.$ (5.8)
Since $\zeta$ is parallel with respect to $\nabla^{1}$, then $D^{1,A}(\varepsilon\cdot\zeta)=(D^{1,A}\varepsilon)\cdot\zeta,$ in consequence,
$(D^{1,A}\varepsilon)\cdot\zeta=-\frac{\tau_{0}}{2}\,\varepsilon\cdot\zeta+2\,(\varepsilon\;\lrcorner\;\tau_{27})\cdot\zeta.$ (5.9)
Hence, the infinitesimal deformation space of a $\mathrm{G}_{2}$-instanton $A$ on a $7$-manifold with coclosed $\mathrm{G}_{2}$-structure is isomorphic to the kernel of the operator
$D^{1,A}+\frac{1}{2}\,\tau_{0}\,\mathrm{Id}_{\Omega^{1}(M,\mathrm{Ad}(P))}-2\,\tau_{27}^{*}\otimes\mathrm{Id}_{\mathrm{Ad}(P)}\in\mathrm{End}\bigl(\Omega^{1}(M,\mathrm{Ad}(P))\bigr).$ (5.10)
We have $D^{t,A}=D^{1,A}+(t-1)\,\frac{1}{4}\,g^{-1}\not${H}. Consequently,
$D^{t,A}(\varepsilon\cdot\zeta)=D^{1,A}(\varepsilon\cdot\zeta)+(t-1)\,\frac{\tau_{0}}{8}\,\varepsilon\cdot\zeta+(t-1)\,3\,(\varepsilon\;\lrcorner\;\tau_{27})\cdot\zeta.$
Setting $t=\frac{1}{3}$, we obtain $D^{1/3,A}(\varepsilon\cdot\zeta)=D^{1,A}(\varepsilon\cdot\zeta)-\frac{\tau_{0}}{12}\,\varepsilon\cdot\zeta-2\,(\varepsilon\;\lrcorner\;\tau_{27})\cdot\zeta$ and consequently,
$D^{1/3,A}+\frac{7}{12}\,\tau_{0}\,\mathrm{Id}=D^{1,A}+\frac{1}{2}\,\tau_{0}\,\mathrm{Id}-2\,\tau_{27}^{*}\otimes\mathrm{Id}_{\mathrm{Ad}(P)}$ (5.11)
5.2 Weitzenböck inequality and curvature obstruction
It is known the Weitzenböck–Lichnerowicz formula with torsion cf. [Agricola2004Holonomy]
Proposition 5.6.
[…]
$(D^{t/3,A})^{2}\mu=(\nabla^{t,A})^{\ast}\nabla^{t,A}\mu+\frac{1}{4}\,\mathrm{Scal}\,\mu+\frac{t}{6}\,d\varphi\cdot\mu-\frac{t^{2}}{18}\,\|\varphi\|^{2}\mu+F\cdot\mu.$ (5.12)
As an application of Propositions 5.5 and 5.6, we prove the following rigidity result.
Corollary 5.7.
Let $(V^{5,2},\varphi)$ be the Stiefel manifold endowed with the coclosed $\mathrm{G}_{2}$–structure given in Corollary 2.8 and let AA denote the $\mathrm{G}_{2}$–instanton (4.9) on the ${\rm SO}(3)$-bundle $P_{1}=V^{5,2}\times_{\phi_{1}}{\rm SO}(3)$. Then, for $y_{2}>0$ as in (4.12) there exists a non–empty interval $y_{2}\in I$ such that, for all values $y_{2}\in I$, the $\mathrm{G}_{2}$–instanton $A$ is rigid for these values of the metric parameter $y_{2}$.
Proof.
[…]
$\bigl|\nabla^{1/3,A}(\varepsilon\cdot\zeta)\bigr|^{2}$ $=\left\langle(D^{\frac{1}{3},A})^{2}(\varepsilon\cdot\zeta),\,\varepsilon\cdot\zeta\right\rangle-\frac{1}{4}\,\mathrm{Scal}\,|\varepsilon\cdot\zeta|^{2}-\frac{1}{18}\,\langle d\varphi\cdot(\varepsilon\cdot\zeta),\,\varepsilon\cdot\zeta\rangle$ (5.13)
$\quad+\frac{1}{162}\,\|\varphi\|^{2}|\varepsilon\cdot\zeta|^{2}-\langle F_{A}(\varepsilon\cdot\zeta),\,\varepsilon\cdot\zeta\rangle.$
From Proposition 5.5, if $\alpha\in\Omega^{1}(M,\mathrm{Ad}(P))$ is an infinitesimal deformation of the $\mathrm{G}_{2}$–instanton $A$, then $\varepsilon$ satisfies $D^{1/3,A}(\varepsilon\cdot\zeta)=-\frac{7}{12}\,\tau_{0}\,\varepsilon\cdot\zeta,$ hence
$\left\langle(D^{\frac{1}{3},A})^{2}(\varepsilon\cdot\zeta),\,\varepsilon\cdot\zeta\right\rangle=\frac{49}{144}\,\tau_{0}^{2}\,|\varepsilon\cdot\zeta|^{2}.$ (5.14)
From (2.23), we compute
$d\varphi$ $=2\left(X^{2356}+X^{2457}+X^{3467}\right)-\frac{3P_{1}}{2y_{2}}\left(X^{1237}-X^{1246}+X^{1345}-X^{1567}\right)$
$+\frac{3P_{3}}{2y_{2}}\left(X^{1234}-X^{1267}+X^{1357}-X^{1456}\right).$
where $P_{1}=a(a^{2}+b^{2})^{-1/2}$ and $P_{3}=b(a^{2}+b^{2})^{-1/2}$. Notice that $P_{1}^{2}+P_{3}^{2}=1$ and $P_{1}+P_{2}\leq 1$. On the other hand, from the proof of Theorem 2.1, we have $\tau_{0}=\frac{6}{7}\left(2+\frac{1}{y_{2}}\right)$ and therefore,
$\ast\tau_{3}$ $=d\varphi-\tau_{0}\psi$
$=$ $\left(\frac{26}{7}+\frac{6}{7y_{2}}\right)\left(X^{2356}+X^{2456}+X^{3467}\right)-\left(\frac{6}{7}+\frac{33}{14}\frac{P_{1}}{y_{2}}\right)\left(X^{1237}-X^{1246}+X^{1345}-X^{1567}\right)$
$+\left(\frac{6}{7}+\frac{33}{14}\frac{P_{3}}{y_{2}}\right)\left(X^{1234}-X^{1267}+X^{1357}-X^{1456}\right).$
Hence $\lvert\tau_{3}\rvert^{2}=3\left(\frac{26}{7}+\frac{6}{7y_{2}}\right)^{2}+4\left(\frac{6}{7}+\frac{33}{14}\frac{P_{1}}{y_{2}}\right)^{2}+4\left(\frac{6}{7}+\frac{33}{14}\frac{P_{3}}{y_{2}}\right)^{2}$ and using $P_{1}+P_{3}\leq P_{1}^{2}+P_{3}^{2}=1,$ we get
$\lvert\tau_{3}\rvert^{2}\leq\frac{4}{49}\left(241+276\,\frac{1}{y_{2}}+\frac{1125}{4}\,\frac{1}{y_{2}^{2}}\right).$ (5.15)
From (5.13), using (5.14)
$(*)$ $\leq\left(\frac{49\lvert\tau_{0}\rvert^{2}}{144}-\frac{\mathrm{Scal}}{4}-\frac{\lvert\tau_{0}\rvert}{18}+\frac{7}{162}+\frac{1}{18}\lvert\tau_{3}\rvert+\lvert F_{A}\rvert\right)\lvert\varepsilon\cdot\zeta\rvert^{2}.$
Thus, from (4.12) and (5.15), we obtain
$0$ $\leq\left(\frac{49\lvert\tau_{0}\rvert^{2}}{144}-\frac{\mathrm{Scal}}{4}-\frac{\lvert\tau_{0}\rvert}{8}+\frac{7}{162}+\frac{\lvert\tau_{3}\rvert}{18}+\lvert F_{A}\rvert\right)\lvert\varepsilon\rvert^{2}$ (5.16)
$\leq\Bigg[\frac{1}{4}\left(2+\frac{1}{y_{2}}\right)^{2}+3-\frac{6}{y_{2}}-\frac{3}{28}\left(2+\frac{1}{y_{2}}\right)+\frac{7}{162}+\frac{1}{63}\sqrt{241+\frac{276}{y_{2}}+\frac{1125}{4y_{2}^{2}}}+\frac{\sqrt{3}}{y_{2}}\Bigg]\lvert\varepsilon\rvert^{2}.$
Setting $x=\frac{1}{y_{2}},$ the right–hand side of (5.16) can be written as $f(x)\,|\varepsilon|^{2},$ where for $x\geq 0$
$f(x)$= $\frac{1}{4}(2+x)^{2}+3-6x-\frac{3}{28}(2+x)+\frac{7}{162}+\frac{1}{63}\sqrt{241+276x+\frac{1125}{4}x^{2}}+\sqrt{3}\,x$
The function $f$ is continuous on $0,\infty)$, moreover, a direct evaluation shows that $f(x)$ becomes negative on a bounded interval. In particular, $f(x)<0\quad\text{for all }x\in(\frac{5}{3},10).$ [^5] This interval is not maximal Equivalently, $f(y_{2})<0$ for all $y_{2}\in\left(\tfrac{1}{10},\tfrac{3}{5}\right)$. For such values of the metric parameter $y_{2}$, the right–hand side of the Weitzenböck inequality becomes strictly negative (see Figure 1), while the left–hand side is non–negative. This yields a contradiction unless $\varepsilon\equiv 0.$ Therefore, the instanton is rigid for this range of metric parameters.
xxf(x)f(x)
Figure 1: Plot of the right–hand side of the Weitzenböck inequality as a function of the metric parameter $y_{2}$. This shows that the expression becomes strictly negative for a non-empty interval of positive values of $y_{2}$.